### v2

# The $D$-equivalence conjecture for hyper-Kähler varieties via hyperholomorphic bundles

Abstract.
We show that birational hyper-Kähler varieties of $K3^{[n]}$-type are derived equivalent, establishing the $D$-equivalence conjecture in these cases. The Fourier–Mukai kernels of our derived equivalences are constructed from projectively hyperholomorphic bundles, following ideas of Markman.
0. Introduction
Throughout, we work over the complex numbers ${\mathbb{C}}$. We recall that the $D$-equivalence conjecture (c.f. [1, 13]) predicts that birational Calabi–Yau varieties have equivalent bounded derived category of coherent sheaves.
Conjecture 0.1 ($D$-equivalence conjecture).
If $X,X^{\prime}$ are nonsingular projective birational Calabi–Yau varieties, then there is an equivalence of bounded derived categories
$D^{b}(X)\simeq D^{b}(X^{\prime}).$
The purpose of this paper is to prove Conjecture 0.1 for hyper-Kähler varieties of $K3^{[n]}$-type; these are nonsingular projective varieties deformation equivalent to the Hilbert scheme of $n$ points on a $K3$ surface.
Theorem 0.2.
Conjecture 0.1 holds for any hyper-Kähler varieties of $K3^{[n]}$-type.
The $D$-equivalence conjecture has been proven by Bridgeland [2] for Calabi–Yau threefolds. For projective hyper-Kähler fourfolds, the $D$-equivalence conjecture holds by combining the classification results [4, 20] and Namikawa’s result for Mukai flops [17]. However, very few cases of this conjecture are known in dimension $>4$. Using equivalences obtained from window conditions, Halpern-Leistner [8] proved the $D$-equivalence conjecture for any hyper-Kähler variety which can be realized as a Bridgeland moduli space of stable objects on a (possibly twisted) $K3$ surface. Theorem 0.2 generalizes Halpern–Leistner’s result, but our approach for constructing the derived equivalences is very different. We obtain explicit Fourier–Mukai kernels which rely on the theory of moduli spaces of hyper-Kähler manifolds and hyperholomorphic bundles [19, 15]; it is also closer in spirit to the proposal of [11, Section 5.1]. It would be interesting to find connections between the two methods.
Acknowledgements
We are grateful to Daniel Huybrechts, Zhiyuan Li, and Alex Perry for helpful discussions. D.M. was supported by a Simons Investigator Grant. J.S. was supported by the NSF grant DMS-2301474 and a Sloan Research Fellowship.
1. Moduli of Hodge isometries
Assume $n\geq 2$. We denote by $\Lambda$ the $K3^{[n]}$-lattice, which is isometric to $H^{2}(X,{\mathbb{Z}})$ for any hyper-Kähler manifold $X$ of $K3^{[n]}$-type. [^1] When we say that $X$ is a hyper-Kähler manifold or a manifold of $K3^{[n]}$-type, it means that $X$ is not necessarily projective. A marking $(X,\eta_{X})$ for a manifold $X$ of $K3^{[n]}$-type is an isometry $\eta_{X}:H^{2}(X,{\mathbb{Z}})\xrightarrow{\simeq}\Lambda$.
1.1. Moduli spaces
We first recall the moduli space of Hodge isometries; this was used by Buskin and Markman to construct algebraic cycles realizing (rational) Hodge isometries [5, 15].
For $\phi\in O(\Lambda_{\mathbb{Q}})$, we define ${\mathfrak{M}}_{\phi}$ to be the moduli space of isomorphism classes of quadruples $(X,\eta_{X},Y,\eta_{Y})$ where $(X,\eta_{X}),(Y,\eta_{Y})$ are the corresponding markings and
$\eta_{Y}^{-1}\circ\phi\circ\eta_{X}:H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$
is a Hodge isometry sending some Kähler class to a Kähler class. The space ${\mathfrak{M}}_{\phi}$ is naturally a non-Hausdorff complex manifold. Let ${\mathfrak{M}}_{\Lambda}$ be the moduli space of marked $K3^{[n]}$-type manifolds $(X,\eta_{X})$. Then we have the natural forgetful maps
$\Pi_{1}:{\mathfrak{M}}_{\phi}\to{\mathfrak{M}}_{\Lambda},\quad(X,\eta_{X},Y,\eta_{Y})\mapsto(X,\eta_{X}),$
$\Pi_{2}:{\mathfrak{M}}_{\phi}\to{\mathfrak{M}}_{\Lambda},\quad(X,\eta_{X},Y,\eta_{Y})\mapsto(Y,\eta_{Y}).$
Any connected component ${\mathfrak{M}}^{0}_{\phi}$ of ${\mathfrak{M}}_{\phi}$ maps to a connected component of ${\mathfrak{M}}_{\Lambda}$ via $\Pi_{i}$ which we denote by ${\mathfrak{M}}^{0}_{\Lambda}$.
Lemma 1.1 ([15, Lemma 5.7]).
The maps $\Pi_{i}:{\mathfrak{M}}^{0}_{\phi}\to{\mathfrak{M}}^{0}_{\Lambda}$ ($i=1,2$) between connected components are surjective.
Now assume that $(X,\eta_{X})\in{\mathfrak{M}}^{0}_{\Lambda}$ and $X^{\prime}$ is a hyper-Kähler manifold bimeromorphic to $X$. Then there is a canonical identification $H^{2}(X,{\mathbb{Z}})=H^{2}(X^{\prime},{\mathbb{Z}})$ and the induced marking $(X^{\prime},\eta_{X^{\prime}})$ lies in the same component ${\mathfrak{M}}^{0}_{\Lambda}$ such that $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ are inseparable points [9].
Lemma 1.2.
Assume that the point $(X,\eta_{X},Y,\eta_{Y})$ lies in a connected component ${\mathfrak{M}}^{0}_{\phi}$. Assume further that $X,X^{\prime}$ are bimeromorphic hyper-Kähler manifolds with the marking $(X^{\prime},\eta_{X^{\prime}})$ induced by $(X,\eta_{X})$. Then the point $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})$ lies in the same connected component ${\mathfrak{M}}^{0}_{\phi}$.
Proof.
Both $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$ which we call $\mathfrak{M}^{0}_{\Lambda}$. We first find paths in ${\mathfrak{M}}^{0}_{\Lambda}$ connecting both points to $(X_{0},\eta_{X_{0}})\in{\mathfrak{M}}^{0}_{\Lambda}$ with $\mathrm{Pic}(X_{0})=0$. By Lemma 1.1, we can lift these paths to ${\mathfrak{M}}^{0}_{\phi}$, which connect $(X,\eta_{X},Y,\eta_{Y})$ to $(X_{0},\eta_{X_{0}},Y_{0},\eta_{Y_{0}})$, and $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})$ to $(X_{0},\eta_{X_{0}},Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$ respectively. On one hand, by considering the projection $\Pi_{2}$, we know that the two points $(Y_{0},\eta_{Y_{0}}),(Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$ lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$; on the other hand, the Hodge isometry condition ensures that both of them share the same period [15, Lemma 5.11] and they have trivial Picard group. By the global Torelli theorem, we must have $(Y_{0},\eta_{Y_{0}})=(Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$. This completes the proof. ∎
Suppose we are given a quadruple $(X,\eta_{X},Y,\eta_{Y})$ in ${\mathfrak{M}}_{\phi}$, and Kähler classes $\omega_{X},\omega_{Y}$ on $X,Y$ which are identified via $\eta_{Y}^{-1}\circ\phi\circ\eta_{X}$. Using this data, one can define a diagonal twistor line $\ell\subset{{\mathfrak{M}}}_{\phi}$ which lifts the twistor lines associated to $(X,\omega_{X})$ and $(Y,\omega_{Y})$ on $\mathfrak{M}_{\Lambda}$. A generic diagonal twistor path on ${\mathfrak{M}}_{\phi}$ is given by a chain of diagonal twistor lines such that, at each node of the chain, the associated hyper-Kähler manifolds have trivial Picard group. Generic diagonal twistor paths are used in Theorem 1.4 below to deform certain Fourier–Mukai kernels.
1.2. Brauer groups
Assume that $X$ is a manifold of $K3^{[n]}$-type. Since $X$ has no odd cohomology, the discussion in [7, Section 4.1] yields the following explicit description of the (cohomological) Brauer group:
(1) $\mathrm{Br}(X)=\left(H^{2}(X,{\mathbb{Z}})/\mathrm{Pic}(X)\right)\otimes{\mathbb{Q}}/{\mathbb{Z}}.$
In particular, for bimeromorphic hyper-Kähler manifolds $X,X^{\prime}$, we have a canonical identification
$\mathrm{Br}(X)=\mathrm{Br}(X^{\prime})$
since both $H^{2}(-,{\mathbb{Z}})$ and $\mathrm{Pic}(-)$ are identified for $X$ and $X^{\prime}$. The description (1) also allows us to present a Brauer class in the form
(2) $\left[\frac{\beta}{m}\right]\in\mathrm{Br}(X),\quad\beta\in H^{2}(X,{\mathbb{Z}}),\quad m\in{\mathbb{N}};$
this is now referred to as the “$B$-field”. The next lemma follows from (1).
Lemma 1.3.
The class (2) vanishes if and only if
$\beta=v+m\beta_{0}\in H^{2}(X,{\mathbb{Z}}),\quad v\in\mathrm{Pic}(X),\quad\beta_{0}\in H^{2}(X,{\mathbb{Z}}).$
We note that the cohomology $H^{2}(X,{\mathbb{Z}})$ forms a trivial local system over any connected component of the moduli space ${\mathfrak{M}}^{0}_{\Lambda}$; therefore (2) for a single $X$ presents a Brauer class for any point in the component ${\mathfrak{M}}^{0}_{\Lambda}$ containing $(X,\eta_{X})$.
1.3. Projectively hyperholomorphic bundles
[…]
$\mathrm{rk}\left({\mathcal{U}}^{[n]}\right)=n!r^{n};$
see [15, Lemma 7.1]. This vector bundle induces a derived equivalence
(3) $\Phi_{{\mathcal{U}}^{[n]}}:D^{b}(M^{[n]})\xrightarrow{\simeq}D^{b}(S^{[n]}).$
Markman further showed in [15, Section 5.6] that the characteristic class of ${\mathcal{U}}^{[n]}$ induces a Hodge isometry
$\phi_{{\mathcal{U}}^{[n]}}:H^{2}(M^{[n]},{\mathbb{Q}})\to H^{2}(S^{[n]},{\mathbb{Q}}).$
The key results, which are summarized in the following theorem, show that the Fourier–Mukai kernel ${\mathcal{U}}^{[n]}$, as a (possibly twisted) projectively hyperholomorphic bundle, deforms along generic diagonal twistor paths. Moreover, at each point of the path, it induces a derived equivalence:
Theorem 1.4 ([15, 12]).
There exist markings $\eta_{M^{[n]}},\eta_{S^{[n]}}$ for the Hilbert schemes $M^{[n]},S^{[n]}$ respectively, which induce $\tau\in O(\Lambda_{\mathbb{Q}})$ via $\phi_{{\mathcal{U}}^{[n]}}$, such that the connected component containing the quadruple
$(M^{[n]},\eta_{M^{[n]}},S^{[n]},\eta_{S^{[n]}})\in{\mathfrak{M}}^{0}_{\tau}$
satisfies the following:
  1. (a)
    For every point $(X,\eta_{X},Y,\eta_{Y})$ lying in the component ${\mathfrak{M}}^{0}_{\tau}$, there exists a twisted vector bundle $({\mathcal{E}},\alpha_{\mathcal{E}})$ over $X\times Y$, which is deformed from ${\mathcal{U}}^{[n]}$ along a generic diagonal twistor path.
  2. (b)
    Using the form (2), the Brauer class in (a) is presented by
    $\alpha_{\mathcal{E}}=\left[-\frac{c_{1}({\mathcal{U}}^{[n]})}{n!r^{n}}\right].$
    Here we view $H^{2}(X\times Y,{\mathbb{Z}})=H^{2}(X,{\mathbb{Z}})\oplus H^{2}(Y,{\mathbb{Z}})$ as a trivial local system over the moduli space ${\mathfrak{M}}^{0}_{\tau}$ via the markings.
  3. (c)
    Further assume that $X,Y$ are varieties. Then the twisted bundle $({\mathcal{E}},\alpha_{\mathcal{E}})$ induces an equivalence of twisted derived categories
    $\Phi_{({\mathcal{E}},\alpha_{\mathcal{E}})}:D^{b}(X,\alpha_{X})\xrightarrow{\simeq}D^{b}(Y,\alpha_{Y}),\quad\alpha_{X}=\left[\frac{a_{X}}{n!r^{n}}\right],\quad\alpha_{Y}=\left[-\frac{a_{Y}}{n!r^{n}}\right],$
    where $a_{X}\in H^{2}(X,{\mathbb{Z}}),a_{Y}\in H^{2}(Y,{\mathbb{Z}})$ are given by
    $c_{1}({\mathcal{U}}^{[n]})=a_{X}+a_{Y}\in H^{2}(X,{\mathbb{Z}})\oplus H^{2}(Y,{\mathbb{Z}}).$
Proof.
(a) was proven in [15, Theorem 8.4]; Markman showed that ${\mathcal{U}}^{[n]}$ over $M^{[n]}\times S^{[n]}$ is projectively slope-stable hyperholomorphic in the sense of [19, 14] which allows him to deform it along diagonal twistor paths to all points in the component ${\mathfrak{M}}^{0}_{\tau}$.
(b) can be obtained by applying Căldăraru’s result [7, Theorem 4.1] along the diagonal twistor paths; see the discussion in [12, Section 2.3].
(c) was proven in [12, Theorem 2.3]. More precisely, the condition that a twisted bundle induces a derived equivalence can be characterized by cohomological properties [6, Theorem 3.2.1]. These properties are preserved along a twistor path due to the fact that cohomology of slope-polystable hyperholomorphic bundles are invariant under hyper-Kähler rotations [18, Corollary 8.1]. Therefore we ultimately reduce the desired cohomological properties to those for $M^{[n]}\times S^{[n]}$ which are given by the original equivalence (3). ∎
1.4. Strategy of the proof
We discuss the strategy of the proof of Theorem 0.2.
We start with a choice of $S$ and $v_{0}=(r,mH,s)$ as in the beginning of Section 1.3, which yields the isometry $\phi_{{\mathcal{U}}^{[n]}}$. Associated to these, we have the moduli space of Hodge isometries ${\mathfrak{M}}_{\tau}$, and the component ${\mathfrak{M}}^{0}_{\tau}$ that contains the quadruple $(M^{[n]},\eta_{M^{[n]}},S^{[n]},\eta_{S^{[n]}})$.
Now assume that $X$ is a variety of $K3^{[n]}$-type. By Lemma 1.1, we can complete it to a quadruple [^3] Here the orders of $X,Y$ are switched as we would like $X$ to be deformed from $S^{[n]}$ later in Section 2.
$(Y,\eta_{Y},X,\eta_{X})\in{\mathfrak{M}}^{0}_{\tau}.$
By Theorem 1.4(b, c), we obtain Brauer classes $\alpha_{X},\alpha_{Y}$ on $X,Y$ respectively, such that
(4) $D^{b}(Y,\alpha_{Y})\simeq D^{b}(X,\alpha_{X}).$
Here the Brauer classes $\alpha_{X},\alpha_{Y}$ only depend on the markings $(X,\eta_{X}),(Y,\eta_{Y})$ respectively.
Let $X^{\prime}$ be another variety of $K3^{[n]}$-type which is birational to $X$. Then the marking $(X,\eta_{X})$ induces a marking $(X^{\prime},\eta_{X^{\prime}})$. By Lemma 1.2 we have
$(Y,\eta_{Y},X^{\prime},\eta_{X}^{\prime})\in{\mathfrak{M}}^{0}_{\tau},$
which yields
(5) $D^{b}(Y,\alpha_{Y})\simeq D^{b}(X^{\prime},\alpha_{X^{\prime}}).$
Combining (4) and (5) yields a proof of a twisted version of Theorem 0.2
(6) $D^{b}(X,\alpha_{X})\simeq D^{b}(X^{\prime},\alpha_{X^{\prime}})$
whose Fourier–Mukai kernel is given by the composition of two (twisted) hyperholomorphic bundles.
In the next section, we show that for any $K3^{[n]}$-type variety $X$, a careful choice of the $K3$ surface $S$ and the Mukai vector $v_{0}=(r,mH,s)$ as in Section 1.3 can make the induced Brauer class $\alpha_{X}$ vanish. In particular, we have
$\alpha_{X}=\alpha_{X^{\prime}}=0,$
and (6) recovers Theorem 0.2.
2. Vanishing of Brauer classes
From now on, we work with a primitively polarized $K3^{[n]}$-type variety $(X,D)$.
2.1. Hilbert schemes
Let $r,s$ be coprime integers with $r\geq 2,s\geq 1$. Let $S$ be a very general projective $K3$ surface with
$\mathrm{Pic}(S)={\mathbb{Z}}H,\quad(mH)^{2}=2rs.$
As in the beginning of Section 1.3, we define $M$ to be the moduli space of stable vector bundles on $S$ with respect to the Mukai vector $v_{0}=(r,mH,s)$. A universal bundle ${\mathcal{U}}$ on $M\times S$ satisfies
$c_{1}({\mathcal{U}})=a_{M}+mH\in H^{2}(M,{\mathbb{Z}})\oplus H^{2}(S,{\mathbb{Z}}),\quad a_{M}\in H^{2}(M,{\mathbb{Z}}),\quad H\in H^{2}(S,{\mathbb{Z}}).$
We denote by $\delta\in\mathrm{Pic}(S^{[n]})$ the half of the exceptional divisor, and we have a canonical isomorphism
(7) $H^{2}(S^{[n]},{\mathbb{Z}})=H^{2}(S,{\mathbb{Z}})\oplus{\mathbb{Z}}\delta.$
We write
$c_{1}({\mathcal{U}}^{[n]})=a_{M^{[n]}}+a_{S^{[n]}}\in H^{2}(M^{[n]},{\mathbb{Z}})\oplus H^{2}(S^{[n]},{\mathbb{Z}})$
with
$a_{M^{[n]}}\in H^{2}(M^{[n]},{\mathbb{Z}}),\quad a_{S^{[n]}}\in H^{2}(S^{[n]},{\mathbb{Z}}).$
By [15, Equation (7.11)], we can present the class $a_{S^{[n]}}$ using (7):
$a_{S^{[n]}}=n!r^{n}\cdot\left(\frac{mH}{r}-\frac{\delta}{2}\right)\in H^{*}(S^{[n]},{\mathbb{Z}}).$
Proposition 2.1.
If there is a $K3$ surface $S$ and a Mukai vector $v_{0}=(r,mH,s)$ as in Section 1.3, and a parallel transport
$\rho:H^{2}(S^{[n]},{\mathbb{Q}})\to H^{2}(X,{\mathbb{Q}})$
satisfying that
(8) $\rho\left(\frac{mH}{r}-\frac{\delta}{2}\right)=\epsilon D+L,\quad\epsilon\in{\mathbb{Q}},\quad L\in H^{2}(X,{\mathbb{Z}}),$
then the $D$-equivalence conjecture holds for $X$.
Proof.
The $K3$ surface $S$ and the Mukai vector $v_{0}$ as above induces the moduli space ${\mathfrak{M}}^{0}_{\tau}$ of Hodge isometries. As in Section 1.4, we can find a point
$(Y,\eta_{Y},X,\eta_{X})\in{\mathfrak{M}}^{0}_{\tau}$
such that the Brauer class $\alpha_{X}$ is given by
$\alpha_{X}=\left[-\frac{\rho(a_{S^{[n]}})}{n!r^{n}}\right]=\left[-\epsilon D-L\right].$
By Lemma 1.3 we have $\alpha_{X}=0$. Hence for any hyper-Kähler variety $X^{\prime}$ birational to $X$, we also have $\alpha_{X^{\prime}}=0$, which induces
$D^{b}(X)\simeq D^{b}(Y,\alpha_{Y})\simeq D^{b}(X^{\prime}).\qed$
2.2. Parallel transports
We construct in this section the $K3$ surface $S$, the Mukai vector $v_{0}=(r,mH,s)$, and the parallel transport $\rho$. This completes the proof of Theorem 0.2 by Proposition 2.1.
We start with a parallel transport
$\rho^{\prime}:H^{2}(X,{\mathbb{Z}})\to H^{2}(S^{\prime[n]},{\mathbb{Z}})$
connecting $X$ and the Hilbert scheme associated with a $K3$ surface $S^{\prime}$ such that
$\rho^{\prime}(D)=uH^{\prime}-v\delta^{\prime},$
$(u,v)=1,\quad u\neq 0,\quad\pm v>1,$
where $H^{\prime}\in H^{2}(S^{\prime},{\mathbb{Z}})$ is a primitive integral vector with
$H^{\prime 2}=2d>0.$
Choose $L^{\prime}\in H^{2}(S^{\prime},{\mathbb{Z}})$ to be an integral vector with
$H^{\prime}\cdot L^{\prime}=1,\quad L^{\prime 2}=2e>0,$
which we fix from now on.
We treat the following two cases separately.
Case 1: $u$ is even. We define $\ell:=(d,v)$ which allows us to write
$d=\ell d^{\prime},\quad v=\ell v^{\prime},\quad(d^{\prime},v^{\prime})=1.$
The coprime condition of $d^{\prime},v^{\prime}$ further allows us to find an integer $k$ satisfying
(9) $\quad(\ell,d^{\prime}+kv^{\prime})=1.$
We define an integral class
$\widetilde{H}^{\prime}:=H^{\prime}+(kv)L^{\prime}\in H^{2}(S^{\prime},{\mathbb{Z}}).$
We can calculate its norm
$\widetilde{H}^{\prime 2}=2\widetilde{d^{\prime}},\quad\widetilde{d^{\prime}}:=d+kv+ek^{2}v^{2}>0.$
In particular, we have
$(\widetilde{d^{\prime}},v)=(d,v)=\ell.$
Note that $\widetilde{H}^{\prime}$ may not be algebraic but we can apply a parallel transport
$\rho^{\prime\prime}:H^{2}(S^{\prime},{\mathbb{Z}})\to H^{2}(S,{\mathbb{Z}})$
moving to a projective $K3$ surface $S$ of Picard rank $1$, such that $H:=\rho^{\prime\prime}(\widetilde{H}^{\prime})$ is algebraic. We claim that the $K3$ surface $S$, the Mukai vector
$v_{0}:=\left(\frac{v^{2}}{\ell},\frac{uv}{2\ell}H,\frac{u^{2}\widetilde{d^{\prime}}}{4\ell}\right),\quad v_{0}^{2}=0,$
and the parallel transport
$\rho:=\left(\rho^{\prime\prime[n]}\circ\rho^{\prime}\right)^{-1}:H^{2}(S^{[n]},{\mathbb{Z}})\to H^{2}(S^{\prime[n]},{\mathbb{Z}})\to H^{2}(X,{\mathbb{Z}})$
work for our purpose.
We first check the condition (8). By definition we have
$\rho\left(\frac{\frac{uv}{2\ell}H}{\frac{v^{2}}{\ell}}-\frac{\delta}{2}\right)=\frac{D}{2v}+\frac{u}{2}\cdot k\rho^{\prime-1}(L^{\prime})$
where the first term is proportional to $D$ and the second term is integral.
Then we check the coprime condition. Note that $\ell$ divides $\widetilde{d^{\prime}}$ and we write $\widetilde{d^{\prime}}=\ell\widetilde{d^{\prime\prime}}$. We obtain that
$\left(\frac{v^{2}}{\ell},\frac{u^{2}\widetilde{d^{\prime}}}{4\ell}\right)=\left(\ell v^{\prime 2},\frac{u^{2}\widetilde{d^{\prime\prime}}}{4}\right)=(\ell,\widetilde{d^{\prime\prime}})=(\ell,d^{\prime}+kv^{\prime})=1$
where we used (9) in the last equation.
Case 2: $u$ is odd. We now define $\ell:=(d,2v)$ and write
$d=\ell d^{\prime},\quad 2v=\ell v^{\prime},\quad(d^{\prime},v^{\prime})=1.$
As before, we find an integer $k$ satisfying
(10) $(\ell,d^{\prime}+kv^{\prime})=1.$
We now define the class
$\widetilde{H}^{\prime}:=H^{\prime}+(2kv)L^{\prime}\in H^{2}(S^{\prime},{\mathbb{Z}})$
which has norm
$\widetilde{H}^{\prime 2}=2\widetilde{d^{\prime}},\quad\widetilde{d^{\prime}}:=d+2kv+4ek^{2}v^{2}>0,$
where we have $(\widetilde{d^{\prime}},2v)=\ell$. We again pass to a projective $K3$ surface $S$ of Picard rank $1$ via a parallel transport operator
$\rho^{\prime\prime}:H^{2}(S^{\prime},{\mathbb{Z}})\to H^{2}(S,{\mathbb{Z}})$
so that $H:=\rho^{\prime\prime}(\widetilde{H}^{\prime})$ is algebraic.
We work with the surface $S$, the parallel transport operator $\rho:=\left(\rho^{\prime\prime[n]}\circ\rho^{\prime}\right)^{-1}$, and now take the Mukai vector
$v_{0}:=\left(\frac{4v^{2}}{\ell},\frac{2uv}{\ell}H,\frac{u^{2}\widetilde{d^{\prime}}}{\ell}\right),\quad v_{0}^{2}=0.$
For condition (8), we have
$\rho\left(\frac{2\frac{uv}{\ell}H}{\frac{4v^{2}}{\ell}}-\frac{\delta}{2}\right)=\frac{D}{2v}+u\cdot k\rho^{\prime-1}(L^{\prime})$
where the first term is proportional to $D$ and the second term is integral.
For the coprime condition, we write $\widetilde{d^{\prime}}=\ell\widetilde{d^{\prime\prime}}$ and have that
$\left(\frac{4v^{2}}{\ell},\frac{u^{2}\widetilde{d^{\prime}}}{\ell}\right)=(\ell v^{\prime 2},u^{2}\widetilde{d^{\prime\prime}})=(\ell,\widetilde{d^{\prime\prime}})=(\ell,d^{\prime}+kv^{\prime})=1$
using (10) in the last step.
This concludes the proof of Theorem 0.2. ∎