### v4

# The $D$-equivalence conjecture for hyper-Kähler varieties via hyperholomorphic bundles

Abstract.
We show that birational hyper-Kähler varieties of $K3^{[n]}$-type are derived equivalent, establishing the $D$-equivalence conjecture in these cases. The Fourier–Mukai kernels of our derived equivalences are constructed from projectively hyperholomorphic bundles, following ideas of Markman. Our method also proves a stronger version of the $D$-equivalence conjecture for hyper-Kähler varieties of $K3^{[n]}$-type with Brauer classes.
0. Introduction
Throughout, we work over the complex numbers ${\mathbb{C}}$. We recall that the $D$-equivalence conjecture [5, 20] predicts that birational Calabi–Yau varieties have equivalent bounded derived categories of coherent sheaves.
Conjecture 0.1 ($D$-equivalence conjecture).
If $X,X^{\prime}$ are nonsingular projective birational Calabi–Yau varieties, then there is an equivalence of bounded derived categories
$D^{b}(X)\simeq D^{b}(X^{\prime}).$
The purpose of this paper is to prove Conjecture 0.1 for hyper-Kähler varieties of $K3^{[n]}$-type; these are nonsingular projective varieties deformation equivalent to the Hilbert scheme of $n$ points on a $K3$ surface. More generally, our method reduces the $D$-equivalence conjecture for hyper-Kähler varieties to the construction of certain projectively hyperholomorphic bundles.
Theorem 0.2.
Conjecture 0.1 holds for any hyper-Kähler varieties of $K3^{[n]}$-type.
The $D$-equivalence conjecture has been proven by Bridgeland [6] for Calabi–Yau threefolds. For projective hyper-Kähler fourfolds, the $D$-equivalence conjecture holds by combining the classification results [8, 29] and the case of Mukai flops by Kawamata [20] and Namikawa [25]. However, very few cases of this conjecture are known in dimension $>4$; see [26, 1] for some partial results. Using equivalences obtained from window conditions, Halpern-Leistner [13] proved the $D$-equivalence conjecture for any hyper-Kähler variety which can be realized as a Bridgeland moduli space of stable objects on a (possibly twisted) $K3$ surface. Theorem 0.2 generalizes Halpern–Leistner’s result, but our construction of the derived equivalences is very different. We obtain explicit Fourier–Mukai kernels which rely on the theory of moduli spaces of hyper-Kähler manifolds and hyperholomorphic bundles [28, 23]; this is closer in spirit to the proposal of Huybrechts [17, Section 5.1]. It would be interesting to find connections between the two approaches.
Our method in fact proves the following stronger, twisted version of the $D$-equivalence conjecture involving arbitrary Brauer classes. Let $X\dashrightarrow X^{\prime}$ be a birational transform between hyper-Kähler varieties of ${K3}^{[n]}$-type. It naturally identifies the Brauer groups of $X,X^{\prime}$: any Brauer class $\alpha\in\mathrm{Br}(X)$ induces a Brauer class $\alpha^{\prime}\in\mathrm{Br}(X^{\prime})$.
Theorem 0.3.
Let $X\dashrightarrow X^{\prime}$ be as above, and let $\alpha$ be any Brauer class on $X$. Then there is an equivalence of bounded derived categories of twisted sheaves
$D^{b}(X,\alpha)\simeq D^{b}(X^{\prime},\alpha^{\prime}).$
Theorem 0.3 specializes to Theorem 0.2 by taking $\alpha=0$.
Acknowledgements
We are grateful to Daniel Huybrechts, Zhiyuan Li, Eyal Markman, Alex Perry, and Ziyu Zhang for helpful discussions. D.M. was supported by a Simons Investigator Grant. J.S. was supported by the NSF grant DMS-2301474 and a Sloan Research Fellowship.
1. Moduli of Hodge isometries
Assume $n\geq 2$. We denote by $\Lambda$ the $K3^{[n]}$-lattice, which is isometric to $H^{2}(X,{\mathbb{Z}})$, equipped with the Beauville–Bogomolov–Fujiki (BBF) form, for any hyper-Kähler manifold $X$ of $K3^{[n]}$-type. [^1] When we say that $X$ is a hyper-Kähler manifold or a manifold of $K3^{[n]}$-type, it means that $X$ is not necessarily projective. In particular, we have a decomposition
$\Lambda=\Lambda_{\mathrm{K3}}\oplus{\mathbb{Z}}\delta,\quad\delta^{2}=2-2n$
with $\Lambda_{\mathrm{K3}}$ the unimodular $K3$ lattice, so that any vector $\omega\in\Lambda$ can be expressed uniquely as
$\omega=\omega_{\mathrm{K3}}+\lambda\delta,\quad\omega_{\mathrm{K3}}\in\Lambda_{\mathrm{K3}},\quad\lambda\in{\mathbb{Z}}.$
A marking $(X,\eta_{X})$ for a manifold $X$ of $K3^{[n]}$-type is an isometry $\eta_{X}:H^{2}(X,{\mathbb{Z}})\xrightarrow{\simeq}\Lambda$.
1.1. Inseparable pairs
We denote by ${\mathfrak{M}}_{\Lambda}$ the moduli space of marked manifolds $(X,\eta_{X})$ of $K3^{[n]}$-type; it is naturally a non-Hausdorff complex manifold whose non-separation illustrates the complexity of the birational/bimeromorphic geometry of hyper-Kähler varieties/manifolds [14].
We say that a pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is inseparable if they represent inseparable points on the moduli space ${\mathfrak{M}}_{\Lambda}$; as a consequence of the global Torelli theorem, this is equivalent to the condition that $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ share the same period and lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$.
Typical examples of inseparable pairs are given by bimeromorphic transforms. More precisely, a bimeromorphic map $X\dashrightarrow X^{\prime}$ induces a natural identification $H^{2}(X,{\mathbb{Z}})=H^{2}(X^{\prime},{\mathbb{Z}})$ respecting the Hodge structures. A marking $\eta_{X}$ for $X$ then induces a marking $\eta_{X^{\prime}}$ for $X^{\prime}$, and the pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is therefore inseparable. Note that inseparable points are not necessarily induced by bimeromorphic transforms directly. As an example, we consider bimeromorphic $X,X^{\prime}$ as above and assume that
$\rho:H^{2}(X^{\prime},{\mathbb{Z}})\to H^{2}(X,{\mathbb{Z}})$
is a parallel transport respecting the Hodge structures. Then the pair
$(X,\eta_{X}),\quad(X^{\prime},\eta_{X^{\prime}}),\quad\eta_{X^{\prime}}:=\eta_{X}\circ\rho$
is inseparable. By [14] (see also [21, Section 3.1]), every inseparable pair arises this way.
1.2. Hodge isometries
We recall the moduli space of Hodge isometries; this was used by Buskin [9] and Markman [23] to construct algebraic cycles realizing rational Hodge isometries.
For $\phi\in O(\Lambda_{\mathbb{Q}})$, we define ${\mathfrak{M}}_{\phi}$ to be the moduli space of isomorphism classes of quadruples $(X,\eta_{X},Y,\eta_{Y})$ where $(X,\eta_{X}),(Y,\eta_{Y})\in\mathfrak{M}_{\Lambda}$ are the corresponding markings and
$\eta_{Y}^{-1}\circ\phi\circ\eta_{X}:H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$
is a Hodge isometry sending some Kähler class of $X$ to a Kähler class of $Y$. We have the natural forgetful maps
$\Pi_{1}:{\mathfrak{M}}_{\phi}\to{\mathfrak{M}}_{\Lambda},\quad(X,\eta_{X},Y,\eta_{Y})\mapsto(X,\eta_{X}),$
$\Pi_{2}:{\mathfrak{M}}_{\phi}\to{\mathfrak{M}}_{\Lambda},\quad(X,\eta_{X},Y,\eta_{Y})\mapsto(Y,\eta_{Y}).$
Any connected component ${\mathfrak{M}}^{0}_{\phi}$ of ${\mathfrak{M}}_{\phi}$ maps to a connected component of ${\mathfrak{M}}_{\Lambda}$ via $\Pi_{i}$ which we denote by ${\mathfrak{M}}^{0}_{\Lambda}$.
Lemma 1.1 ([23, Lemma 5.7]).
The maps $\Pi_{i}:{\mathfrak{M}}^{0}_{\phi}\to{\mathfrak{M}}^{0}_{\Lambda}$ ($i=1,2$) between connected components are surjective.
Lemma 1.2.
Assume that the point $(X,\eta_{X},Y,\eta_{Y})$ lies in a connected component ${\mathfrak{M}}^{0}_{\phi}$. Assume further that $(X,\eta),(X^{\prime},\eta_{X^{\prime}})$ form an inseparable pair such that
(1) $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})\in{\mathfrak{M}}_{\phi}.$
Then $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})$ lies in the same component ${\mathfrak{M}}^{0}_{\phi}$.
Note that (1) is equivalent to the condition that $\phi$ sends some Kähler class of $X^{\prime}$ to a Kähler class of $Y$.
Proof.
Both $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$ which we call $\mathfrak{M}^{0}_{\Lambda}$. We first find paths in ${\mathfrak{M}}^{0}_{\Lambda}$ connecting both points to $(X_{0},\eta_{X_{0}})\in{\mathfrak{M}}^{0}_{\Lambda}$ with $\mathrm{Pic}(X_{0})=0$. By Lemma 1.1, we can lift these paths to ${\mathfrak{M}}_{\phi}$, which connect $(X,\eta_{X},Y,\eta_{Y})$ to $(X_{0},\eta_{X_{0}},Y_{0},\eta_{Y_{0}})$, and $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})$ to $(X_{0},\eta_{X_{0}},Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$ respectively. On one hand, by considering the projection $\Pi_{2}$, we know that the two points $(Y_{0},\eta_{Y_{0}}),(Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$ lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$; on the other hand, the Hodge isometry condition ensures that both of them share the same period [23, Lemma 5.4] and they have trivial Picard group. By the global Torelli theorem, we must have $(Y_{0},\eta_{Y_{0}})=(Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$. This completes the proof. ∎
Suppose we are given a point $(X,\eta_{X},Y,\eta_{Y})$ in ${\mathfrak{M}}_{\phi}$, and Kähler classes $\omega_{X},\omega_{Y}$ on $X,Y$ which are identified via $\eta_{Y}^{-1}\circ\phi\circ\eta_{X}$. Using this data, one can define a diagonal twistor line $\ell\subset{{\mathfrak{M}}}_{\phi}$ which lifts the twistor lines associated to $(X,\omega_{X})$ and $(Y,\omega_{Y})$ on $\mathfrak{M}_{\Lambda}$. A generic diagonal twistor path on ${\mathfrak{M}}_{\phi}$ is given by a chain of diagonal twistor lines such that, at each node of the chain, the associated hyper-Kähler manifolds have trivial Picard group. Generic diagonal twistor paths are used in Theorem 1.3 below to deform certain Fourier–Mukai kernels.
1.3. Brauer groups
Assume that $X$ is a manifold of $K3^{[n]}$-type. Since $X$ has no odd cohomology, the discussion in [11, Section 4.1] yields the following explicit description of the (cohomological) Brauer group:
(2) $\mathrm{Br}(X)=\left(H^{2}(X,{\mathbb{Z}})/\mathrm{Pic}(X)\right)\otimes{\mathbb{Q}}/{\mathbb{Z}}.$
In particular, given a bimeromorphic map $X\dashrightarrow X^{\prime}$ between manifolds of $K3^{[n]}$-type, there is a natural identification
$\mathrm{Br}(X)=\mathrm{Br}(X^{\prime})$
since both $H^{2}(-,{\mathbb{Z}})$ and $\mathrm{Pic}(-)$ are identified for $X$ and $X^{\prime}$. The description (2) also allows us to present a Brauer class in the form
(3) $\left[\frac{\beta}{d}\right]\in\mathrm{Br}(X),\quad\beta\in H^{2}(X,{\mathbb{Z}}),\quad d\in{\mathbb{Z}}_{>0};$
this is referred to as the “$B$-field”.
We note that the cohomology $H^{2}(X,{\mathbb{Z}})$ forms a trivial local system over any connected component of the moduli space ${\mathfrak{M}}^{0}_{\Lambda}$; therefore (3) for a single $X$ presents a Brauer class for any point in the component ${\mathfrak{M}}^{0}_{\Lambda}$ containing $(X,\eta_{X})$.
1.4. Projectively hyperholomorphic bundles
Using the Bridgeland–King–Reid (BKR) correspondence [7], Markman constructed in [23] a class of projectively hyperholomorphic bundles which we recall here. We consider a very general projective $K3$ surface $S$ with $\mathrm{Pic}(S)={\mathbb{Z}}H$. Assume that $r,s$ are two coprime integers with $r\geq 2$. Assume further that the Mukai vector
$v_{0}:=(r,mH,s)\in H^{*}(S,{\mathbb{Z}})$
[…]
$\mathrm{rk}({\mathcal{U}}^{[n]})=n!r^{n};$
see [23, Lemma 7.1]. This vector bundle induces a derived equivalence
(4) $\Phi_{{\mathcal{U}}^{[n]}}:D^{b}(M^{[n]})\xrightarrow{\simeq}D^{b}(S^{[n]}).$
Markman further showed in [23, Section 5.6] that the characteristic class of ${\mathcal{U}}^{[n]}$ induces a Hodge isometry
$\phi_{{\mathcal{U}}^{[n]}}:H^{2}(M^{[n]},{\mathbb{Q}})\to H^{2}(S^{[n]},{\mathbb{Q}}).$
Under the natural identification
(5) $H^{2}(M^{[n]},{\mathbb{Q}})=H^{2}(M,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta,\quad H^{2}(S^{[n]},{\mathbb{Q}})=H^{2}(S,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta,$
this Hodge isometry is of the form
$\phi_{{\mathcal{U}}^{[n]}}=(\phi_{\mathcal{U}},\mathrm{id}),\quad\phi_{\mathcal{U}}:H^{2}(M,{\mathbb{Q}})\to H^{2}(S,{\mathbb{Q}})$
where $\phi_{\mathcal{U}}$ is the Hodge isometry of $K3$ surfaces induced by ${\mathcal{U}}$; see [23, Corollary 7.3].
The key results, which are summarized in the following theorem, show that the Fourier–Mukai kernel ${\mathcal{U}}^{[n]}$, as a projectively hyperholomorphic bundle, deforms along generic diagonal twistor paths. Moreover, at each point of the path, it induces a (twisted) derived equivalence:
Theorem 1.3 ([23, 19]).
There exist markings $\eta_{M^{[n]}},\eta_{S^{[n]}}$ for the Hilbert schemes $M^{[n]},S^{[n]}$ respectively, which induce $\phi\in O(\Lambda_{\mathbb{Q}})$ via $\phi_{{\mathcal{U}}^{[n]}}$, such that the connected component containing the quadruple
$(M^{[n]},\eta_{M^{[n]}},S^{[n]},\eta_{S^{[n]}})\in{\mathfrak{M}}^{0}_{\phi}$
satisfies the following:
  1. (a)
    For every point $(X,\eta_{X},Y,\eta_{Y})$ lying in the component ${\mathfrak{M}}^{0}_{\phi}$, there exists a twisted vector bundle $({\mathcal{E}},\alpha_{\mathcal{E}})$ on $X\times Y$, which is deformed from ${\mathcal{U}}^{[n]}$ along a generic diagonal twistor path.
  2. (b)
    Using the form (3), the Brauer class in (a) is presented by
    $\alpha_{\mathcal{E}}=\left[-\frac{c_{1}({\mathcal{U}}^{[n]})}{\mathrm{rk}({\mathcal{U}}^{[n]})}\right].$
    Here we view $H^{2}(X\times Y,{\mathbb{Z}})=H^{2}(X,{\mathbb{Z}})\oplus H^{2}(Y,{\mathbb{Z}})$ as a trivial local system over the moduli space ${\mathfrak{M}}^{0}_{\phi}$ via the markings.
  3. (c)
    Further assume that $X,Y$ are varieties. Then the twisted bundle $({\mathcal{E}},\alpha_{\mathcal{E}})$ induces an equivalence of twisted derived categories
    $\Phi_{({\mathcal{E}},\alpha_{\mathcal{E}})}:D^{b}(X,\alpha_{X})\xrightarrow{\simeq}D^{b}(Y,\alpha_{Y}),\quad\alpha_{X}=\left[\frac{a_{X}}{\mathrm{rk}({\mathcal{E}})}\right],\quad\alpha_{Y}=\left[-\frac{a_{Y}}{\mathrm{rk}({\mathcal{E}})}\right],$
    where $a_{X}\in H^{2}(X,{\mathbb{Z}}),a_{Y}\in H^{2}(Y,{\mathbb{Z}})$ are given by
    $c_{1}({\mathcal{U}}^{[n]})=a_{X}+a_{Y}\in H^{2}(X,{\mathbb{Z}})\oplus H^{2}(Y,{\mathbb{Z}}).$
Proof.
(a) was proven in [23, Theorem 8.4]; Markman showed that ${\mathcal{U}}^{[n]}$ on $M^{[n]}\times S^{[n]}$ is projectively slope-stable hyperholomorphic in the sense of [28, 22] which allows him to deform it along diagonal twistor paths to all points in the component ${\mathfrak{M}}^{0}_{\phi}$.
(b) can be obtained by applying Căldăraru’s result [11, Theorem 4.1] along the diagonal twistor paths; see the discussion in [19, Section 2.3].
(c) was proven in [19, Theorem 2.3]. More precisely, the condition that a twisted bundle induces a twisted derived equivalence can be characterized by cohomological properties [10, Theorem 3.2.1]. These properties are preserved along a twistor path due to the fact that the cohomology of slope-polystable hyperholomorphic bundles is invariant under hyper-Kähler rotations [27, Corollary 8.1]. Therefore we ultimately reduce the desired cohomological properties to those for $M^{[n]}\times S^{[n]}$ which are given by the original equivalence (4). ∎
1.5. Birational geometry and MBM classes
The birational geometry of hyper-Kähler varieties is governed by certain integral primitive cohomology classes, called the monodromy birationally minimal (MBM) classes. We refer to [3] for an introduction to these classes. In the following, we summarize some results which are needed in our proof.
Let $X$ be a variety of $K3^{[n]}$-type. We consider its birational Kähler cone $\mathcal{BK}_{X}$ and the positive cone $\mathcal{C}_{X}$:
$\mathcal{BK}_{X}\subset\mathcal{C}_{X}\subset H^{1,1}(X,{\mathbb{R}}).$
The positive cone is convex and admits a wall-and-chamber structure. The closure of the birational Kähler cone within the positive cone is a convex sub-cone [15], which inherits a wall-and-chamber structure. Furthermore, by a result of Amerik–Verbitsky [2], and independently Mongardi [24], all the walls are governed by the MBM classes.
Theorem 1.4 ([24, 2]).
Any wall of $\mathcal{C}_{X}$ is described by a hyperplane of the form
${\mathcal{W}}^{\perp}:=\{\omega\in H^{1,1}(X,{\mathbb{R}}),\,(\omega,{\mathcal{W}})=0\}\subset H^{1,1}(X,{\mathbb{R}})$
with ${\mathcal{W}}$ an algebraic MBM class in $\mathrm{Pic}(X)$. Here the pairing is with respect to the BBF form. Moreover, any chamber in $\mathcal{C}_{X}$ can be realized as the Kähler cone of a birational hyper-Kähler $X^{\prime}$ through a parallel transport $\rho:H^{2}(X^{\prime},{\mathbb{Z}})\to H^{2}(X,{\mathbb{Z}})$ respecting the Hodge structures.
Note that any chamber in $\mathcal{BK}_{X}\subset\mathcal{C}_{X}$ is given by the pullback of the Kähler cone via a birational transform $X\dashrightarrow X^{\prime}$ of hyper-Kähler varieties. By the discussions of Section 1.1, any chamber of $\mathcal{C}_{X}$ corresponds to a marked variety $(X^{\prime},\eta_{X^{\prime}})$ of $K3^{[n]}$-type such that the pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is inseparable.
We also need the following boundedness result, which notably implies that wall-and-chamber structure of $\mathcal{C}_{X}$ is locally polyhedral; see [18, Remark 8.2.3] for a proof of the implication. The boundedness result was essentially obtained by [4], as explained in [2, Section 6.2].
Theorem 1.5 ([4, 2]).
There is a constant $C_{0}>0$, such that for any variety $X$ of $K3^{[n]}$-type and any MBM class ${\mathcal{W}}\in H^{2}(X,{\mathbb{Z}})$ we have
$0<-{\mathcal{W}}^{2}<C_{0}.$
Here the norm is with respect to the BBF form.
For any rational Hodge isometry $\phi:H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$ between varieties of $K3^{[n]}$-type, which sends an MBM class ${\mathcal{W}}_{X}$ on $X$ to a class proportional to an MBM class ${\mathcal{W}}_{Y}$ on $Y$, there exist coprime integers $a,b$ such that
$\phi({\mathcal{W}}_{X})=\frac{a}{b}{\mathcal{W}}_{Y}.$
The following is an immediate consequence of Theorem 1.5.
Corollary 1.6.
For any $X,Y,\phi,{\mathcal{W}}_{X},{\mathcal{W}}_{Y}$ as above, we have
$a^{2}<C_{0},\quad b^{2}<C_{0}.$
Proof.
Since $\phi$ is an isometry, we have
$\frac{a^{2}}{b^{2}}=\frac{{\mathcal{W}}_{X}^{2}}{{\mathcal{W}}_{Y}^{2}}$
By Theorem 1.5, both $-{\mathcal{W}}_{X}^{2}$ and $-{\mathcal{W}}_{Y}^{2}$ are positive integers $<C_{0}$. The corollary follows from the assumption that $a,b$ are coprime. ∎
1.6. Proof strategy
We discuss the strategy of the proof of Theorem 0.3; Theorem 0.2 is then deduced as a special case.
Let $X$ be a variety of $K3^{[n]}$-type. It suffices to prove Theorem 0.3 for a hyper-Kähler birational model $X^{\prime}$ with a birational map $X\dashrightarrow X^{\prime}$ which corresponds to a chamber in $\mathcal{BK}_{X}$ adjacent to the Kähler cone of $X$. By Theorem 1.5, the wall between these two chambers is given by an algebraic MBM class ${\mathcal{W}}\in\mathrm{Pic}(X)$.
Now we choose a $K3$ surface $S$ and a Mukai vector $v_{0}=(r,mH,s)$ as in the beginning of Section 1.4, which yields the isometry $\phi_{{\mathcal{U}}^{[n]}}$. Associated to these, we have the moduli space of Hodge isometries ${\mathfrak{M}}_{\phi}$, and the component ${\mathfrak{M}}^{0}_{\phi}$ that contains the quadruple $(M^{[n]},\eta_{M^{[n]}},S^{[n]},\eta_{S^{[n]}})$.
For the given birational $X,X^{\prime}$, by Lemma 1.1, we can complete them to a pair of quadruples
$(Y,\eta_{Y},X,\eta_{X}),\quad(Y^{\prime},\eta_{Y^{\prime}},X^{\prime},\eta_{X^{\prime}})\in{\mathfrak{M}}^{0}_{\phi}$
such that the marking $\eta_{X^{\prime}}$ is induced by $\eta_{X}$ via the birational map $X\dashrightarrow X^{\prime}$. [^3] Here we would like $X,X^{\prime}$ to be deformed from $S^{[n]}$ later in Section 2. In particular, the pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is inseparable. We note that the pair $(Y,\eta_{Y}),(Y^{\prime},\eta_{Y^{\prime}})$ is also inseparable. This is because they share the same period and lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$. Moreover, by definition, $\phi^{-1}$ sends a Kähler class of $X$ (resp. $X^{\prime}$) to a Kähler class of $Y$ (resp. $Y^{\prime}$). [^4] Here we suppress the markings and still use $\phi$ to denote the Hodge isometry $H^{2}(Y,{\mathbb{Q}})\to H^{2}(X,{\mathbb{Q}})$ for notational convenience. Therefore, if
[…]
there must be a point on the wall separating the Kähler cones of $X,X^{\prime}$ which is sent to the interior of a chamber of the positive cone $\mathcal{C}_{Y}$. In particular, there exists a hyper-Kähler birational model ${Y^{\prime\prime}}$ of $Y$ with a marking $({Y^{\prime\prime}},\eta_{Y^{\prime\prime}})$ such that the pair $(Y,\eta_{Y}),(Y^{\prime\prime},\eta_{Y^{\prime\prime}})$ is inseparable and
$(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X,\eta_{X}),\quad(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X^{\prime},\eta_{X^{\prime}})\in{\mathfrak{M}}_{\phi}.$
Furthermore, by Lemma 1.2, both points lie in the connected component we started with:
$(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X,\eta_{X}),\quad(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X^{\prime},\eta_{X^{\prime}})\in{\mathfrak{M}}^{0}_{\phi}.$
By Theorem 1.3(b, c), we obtain Brauer classes $\alpha_{X},\alpha_{Y^{\prime\prime}}$ on $X,Y^{\prime\prime}$ respectively, such that
$D^{b}(Y^{\prime\prime},\alpha_{Y^{\prime\prime}})\simeq D^{b}(X,\alpha_{X}),\quad D^{b}(Y^{\prime\prime},\alpha_{Y^{\prime\prime}})\simeq D^{b}(X^{\prime},\alpha_{X^{\prime}}).$
Here the Brauer classes $\alpha_{X},\alpha_{Y^{\prime\prime}}$ only depend on the markings $(X,\eta_{X}),(Y^{\prime\prime},\eta_{Y^{\prime\prime}})$ respectively, and the Brauer class $\alpha_{X^{\prime}}$ is induced by $\alpha_{X}$. Combining both equations yields
$D^{b}(X,\alpha_{X})\simeq D^{b}(X^{\prime},\alpha_{X^{\prime}})$
whose Fourier–Mukai kernel is the composition of two (twisted) hyperholomorphic bundles.
In the next section, we show that for any pair $X,X^{\prime}$ as above with a Brauer class $\alpha\in\mathrm{Br}(X)$ and an algebraic MBM class ${\mathcal{W}}\in\mathrm{Pic}(X)$, a careful choice of the $K3$ surface $S$ and the Mukai vector $v_{0}=(r,mH,s)$ as in Section 1.4 can simultaneously ensure that the condition (6) holds and the induced Brauer class is as desired:
(7) $\alpha_{X}=\alpha.$
This completes the proof of Theorem 0.3.
Remark 1.7.
For a general birational transform $X\dashrightarrow X^{\prime}$ of varieties of $K3^{[n]}$-type, which do not correspond to adjacent chambers in the birational Kähler cone $\mathcal{BK}_{X}$, our proof realizes the derived equivalence
$D^{b}(X,\alpha)\simeq D^{b}(X^{\prime},\alpha^{\prime})$
via two sequences of varieties $X_{1},\ldots,X_{t-1}$ and $Y_{1},\ldots,Y_{t}$, with each $X_{i}$ birational to $X,X^{\prime}$, such that
(8) $D^{b}(X,\alpha)\simeq D^{b}(Y_{1},\alpha_{Y_{1}})\simeq D^{b}(X_{1},\alpha_{X_{1}})\simeq D^{b}(Y_{2},\alpha_{Y_{2}})\simeq\cdots\\ \simeq D^{b}(Y_{t-1},\alpha_{Y_{t-1}})\simeq D^{b}(X_{t-1},\alpha_{X_{t-1}})\simeq D^{b}(Y_{t},\alpha_{Y_{t}})\simeq D^{b}(X^{\prime},\alpha^{\prime}).$
Each of the derived equivalences in (8) is induced by a (twisted) hyperholomorphic bundle.
2. Proof of Theorem 0.3
From now on, we fix a variety $X$ of $K3^{[n]}$-type, a Brauer class $\alpha\in\mathrm{Br}(X)$, and an algebraic MBM class ${\mathcal{W}}\in\mathrm{Pic}(X)$ as in Section 1.6. In particular, the variety $X$ has Picard rank $\geq 2$. We present the Brauer class $\alpha$ by a class in the rational transcendental lattice $T(X)_{\mathbb{Q}}\subset H^{2}(X,{\mathbb{Q}})$ via (2) and (3):
$\alpha=\left[-\frac{{\mathcal{B}}}{d}\right],\quad{\mathcal{B}}\in T(X),\quad d\in{\mathbb{Z}}_{>0}.$
Up to adjusting $-\frac{{\mathcal{B}}}{d}$ by an integral class in $T(X)\subset H^{2}(X,{\mathbb{Z}})$, we may further assume that the class ${\mathcal{B}}$ satisfies
${\mathcal{B}}^{2}=2e>0.$
2.1. Divisor classes
Recall that the divisibility $\mathrm{div}(\omega)$ of a class $\omega\in H^{2}(X,{\mathbb{Z}})$ is the positive generator of the subgroup
$\{(\omega,\mu),\,\mu\in H^{2}(X,{\mathbb{Z}})\}\subset{\mathbb{Z}}$
Lemma 2.1.
There exists a class ${\mathcal{A}}\in\mathrm{Pic}(X)$ such that
$({\mathcal{A}},{\mathcal{W}})\neq 0,\quad\mathrm{div}({\mathcal{A}})=1.$
Proof.
We pick a marking identifying $H^{2}(X,{\mathbb{Z}})$ with a $K3^{[n]}$-lattice $\Lambda=\Lambda_{\mathrm{K3}}\oplus{\mathbb{Z}}\delta$. For any $g\in O(\Lambda)$, since $g(\delta)^{\perp}$ is a unimodular $K3$-lattice, any primitive vector $\omega\in g(\delta)^{\perp}\subset\Lambda$ satisfies $\mathrm{div}(\omega)=1.$ We would like to choose $g$ so that there exists ${\mathcal{A}}\in g(\delta)^{\perp}\cap\mathrm{Pic}(X)$ satisfying $({\mathcal{A}},{\mathcal{W}})\neq 0$. In other words, we want
$g(\delta)^{\perp}\cap\mathrm{Pic}(X)\neq{\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X).$
If we base change to $\mathbb{C}$, the set of $g\in O(\Lambda)_{\mathbb{C}}$ such that
$g(\delta)^{\perp}\cap\mathrm{Pic}(X)_{\mathbb{C}}\neq{\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X)_{\mathbb{C}}$
is open in the Zariski topology. Furthermore, it is nonempty since $X$ has Picard rank $\geq 2$. Since $O(\Lambda)$ is Zariski-dense in $O(\Lambda)_{\mathbb{C}}$, we can find $g\in O(\Lambda)$ satisfying this condition as well. ∎
Up to replacing ${\mathcal{A}}$ by $-{\mathcal{A}}$, we may assume
$C_{1}:=({\mathcal{A}},{\mathcal{W}})>0$
which we fix from now on.
Proposition 2.2.
For any $N>0$, there exists a class ${\mathcal{D}}\in\mathrm{Pic}(X)$ of divisibility $1$, satisfying
${\mathcal{D}}^{2}>N,\quad({\mathcal{D}},{\mathcal{W}})=C_{1}.$
Proof.
Since $X$ has Picard rank $\geq 2$, we have ${\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X)\neq 0$. Pick an element
$\omega\in{\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X),\quad\omega^{2}>0.$
Then for large enough $t\in{\mathbb{Z}}_{>0}$, we have
$({\mathcal{A}}+t\omega,{\mathcal{W}})=C_{1},\quad({\mathcal{A}}+t\omega)^{2}>N.$
It suffices to show that there exist infinitely many choices of $t\in{\mathbb{Z}}_{>0}$ satisfying
$\mathrm{div}({\mathcal{A}}+t\omega)=1.$
We pick an integral class $\mu\in H^{2}(X,{\mathbb{Z}})$ such that
$({\mathcal{A}},\mu)=1,\quad(\omega,\mu)\neq 0;$
then we pick another integral class $\nu\in H^{2}(X,{\mathbb{Z}})$ such that
$({\mathcal{A}},\nu)=0,\quad(\omega,\nu)\neq 0.$
We claim that for sufficiently large $t\in{\mathbb{Z}}_{>0}$ with $1+t(\omega,\mu)$ a prime number, the class ${\mathcal{A}}+t\omega$ must have divisibility $1$. This follows immediately from the observation that
[…]
2.2. Mukai vectors
We construct the $K3$ surface $S$ and the Mukai vector $v_{0}$ of Section 1.6.
By Proposition 2.2, we can find ${\mathcal{D}}\in\mathrm{Pic}(X)$ with
(9) $\mathrm{div}({\mathcal{D}})=1,\quad({\mathcal{D}},{\mathcal{W}})=C_{1}>0,\quad{\mathcal{D}}^{2}=2g>2C_{0}C_{1}.$
Repeating the same argument as in Proposition 2.2, we also find $t\in{\mathbb{Z}}_{>0}$ such that
$\mathrm{div}({\mathcal{D}}+4gtd{\mathcal{B}})=1.$
By the Eichler criterion (see [12, Lemma 3.5]), there exists a marking
$\eta_{X}:H^{2}(X,{\mathbb{Z}})\xrightarrow{\simeq}\Lambda_{\mathrm{K3}}\oplus{\mathbb{Z}}\delta$
such that
(10) $\eta_{X}({\mathcal{D}}+4gtd{\mathcal{B}})=h-2gtd^{2}\delta$
with $h\in\Lambda_{\mathrm{K3}}$ a primitive vector satisfying
$h^{2}=2g\left(1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)>0.$
Here we have used that $({\mathcal{D}},{\mathcal{B}})=0$ since ${\mathcal{B}}$ is transcendental.
Now let $(S,H)$ be a primitively polarized $K3$ surface of Picard rank $1$ of degree
$H^{2}=2g\left(1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right).$
In view of (10) and up to replacing $\eta_{X}$ by $-\eta_{X}$, there is a parallel transport
$\rho:H^{2}(S^{[n]},{\mathbb{Q}})\to H^{2}(X,{\mathbb{Q}})$
satisfying
(11) $\rho(H-2gtd^{2}\delta)=\epsilon({\mathcal{D}}+4gtd{\mathcal{B}}),$
where $\epsilon=\pm 1$ is a sign determined by the orientation.
We set the Mukai vector
[…]
which clearly satisfies
$\mathrm{gcd}\left(16gt^{2}d^{4},1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)=1,\quad 16gt^{2}d^{4}\geq 2,\quad v_{0}^{2}=0.$
Therefore, the moduli space $M$ of stable vector bundles on $S$ with Mukai vector $v_{0}$ is a $K3$ surface of Picard rank $1$ with a universal bundle ${\mathcal{U}}$ on $M\times S$ which we fix from now on.
Proposition 2.3.
Let $S,M,{\mathcal{U}}$ be as above.
  1. (a)
    The primitive polarization $\widehat{H}$ of $M$ satisfies $\widehat{H}^{2}=H^{2}$.
  2. (b)
    Let $s\in S$ be a point. Assume that the vector bundle ${\mathcal{U}}|_{s}$ has Mukai vector
    $\widehat{v}_{0}=(16gt^{2}d^{4},k\widehat{H},\widehat{s})\in H^{*}(M,{\mathbb{Z}}).$
    Then we have
    $\mathrm{gcd}(16gt^{2}d^{4},k)=4td^{2}.$
Proof.
(a) follows from [30, Appendix A]. For (b), we note that [30, Theorem 2.2] implies that the Mukai vector $\widehat{v}_{0}$ is primitive with $\widehat{v}_{0}^{2}=0$. Using (a), we deduce that
$\widehat{s}=\left(\frac{k}{4td^{2}}\right)^{2}\left(1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)\in{\mathbb{Z}}.$
Therefore, we have that $k$ is divisible by $4td^{2}$, which shows
$4td^{2}\,|\,\mathrm{gcd}(16gt^{2}d^{4},k).$
On the other hand, if $\frac{k}{4td^{2}}$ is not coprime to $16gt^{2}d^{4}$, the Mukai vector $\widehat{v}_{0}$ is divisible by their common factor. This contradicts the fact that $\widehat{v}_{0}$ is primitive. ∎
2.3. End of proof
We complete the proof using the $K3$ surface $S$, the Mukai vector $v_{0}$, and the universal bundle ${\mathcal{U}}$ constructed in the last section. This gives the vector bundle ${\mathcal{U}}^{[n]}$ on $M^{[n]}\times S^{[n]}$. We write
$c_{1}({\mathcal{U}}^{[n]})=a_{M^{[n]}}+a_{S^{[n]}}\in H^{2}(M^{[n]},{\mathbb{Z}})\oplus H^{2}(S^{[n]},{\mathbb{Z}})$
with
$a_{M^{[n]}}\in H^{2}(M^{[n]},{\mathbb{Z}}),\quad a_{S^{[n]}}\in H^{2}(S^{[n]},{\mathbb{Z}}).$
Recall the natural identification
(12) $H^{2}(S^{[n]},{\mathbb{Z}})=H^{2}(S,{\mathbb{Z}})\oplus{\mathbb{Z}}\delta.$
By [23, Equation (7.11)], we can present the class $a_{S^{[n]}}$ using (12):
$a_{S^{[n]}}=\mathrm{rk}({\mathcal{U}}^{[n]})\cdot\left(\frac{\epsilon\cdot 4td^{2}H}{16gt^{2}d^{4}}-\frac{\delta}{2}\right)\in H^{2}(S^{[n]},{\mathbb{Z}}).$
Via the parallel transport (11), we obtain
$\rho\left(\frac{a_{S^{[n]}}}{\mathrm{rk}({\mathcal{U}}^{[n]})}\right)$ $=\rho\left(\frac{\epsilon\cdot H}{4gtd^{2}}-\frac{\delta}{2}\right)$
$=\rho\left(\epsilon\left(\frac{H}{4gtd^{2}}-\frac{\delta}{2}\right)+(\epsilon-1)\frac{\delta}{2}\right)$
$=\frac{\epsilon\cdot\rho(H-2gtd^{2}\delta)}{4gtd^{2}}+\frac{(\epsilon-1)}{2}\rho(\delta)$
$=\frac{{\mathcal{D}}+4gtd{\mathcal{B}}}{4gtd^{2}}+\frac{(\epsilon-1)}{2}\rho(\delta)$
$=\frac{{\mathcal{B}}}{d}+\textup{[class in $\mathrm{Pic}(X)_{\mathbb{Q}}$]}+\textup{[class in $H^{2}(X,{\mathbb{Z}})$]}\in H^{2}(X,{\mathbb{Q}}).$
Hence, by Theorem 1.3(c), we have
$\alpha_{X}=\left[-\rho\left(\frac{a_{S^{[n]}}}{\mathrm{rk}({\mathcal{U}}^{[n]})}\right)\right]=\left[-\frac{{\mathcal{B}}}{d}\right]=\alpha.$
To complete the proof, it remains to address (6). This is given by the following proposition.
Proposition 2.4.
Let ${\mathfrak{M}}_{\phi}^{0}$ be the connected component of the moduli space of Hodge isometries constructed from $S,M,{\mathcal{U}}$ as above. For any quadruple
$(Y,\eta_{Y},X,\eta_{X})\in{\mathfrak{M}}_{\phi}^{0}$
with $X,{\mathcal{W}}$ fixed as above, the class $\phi^{-1}({\mathcal{W}})\in H^{2}(Y,{\mathbb{Q}})$ is not proportional to any MBM class on $Y$. Here we suppress the markings and still use $\phi$ to denote the Hodge isometry $H^{2}(Y,{\mathbb{Q}})\to H^{2}(X,{\mathbb{Q}})$ for notational convenience.
Proof.
The main idea of the argument is that, for our choice of the Mukai vector $v_{0}$, by a calculation of Buskin [9], the rational Hodge isometry $\phi^{-1}$ is conjugate to a reflection by a vector of large norm. By Corollary 1.6, we then show that it cannot send ${\mathcal{W}}$ to a class proportional to an MBM class.
The details are as follows. Since the MBM classes are deformation invariant, we only need to treat the Hodge isometry
$\phi_{{\mathcal{U}}^{[n]}}:H^{2}(M^{[n]},{\mathbb{Q}})\to H^{2}(S^{[n]},{\mathbb{Q}})$
which can be further simplified under the identification (5):
$(\phi_{\mathcal{U}},\mathrm{id}):H^{2}(M,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta\to H^{2}(S,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta.$
Assume that
(13) $\phi_{{\mathcal{U}}^{[n]}}^{-1}(\rho^{-1}({\mathcal{W}}))=\frac{b}{a}{\mathcal{W}}^{\prime}$
with ${\mathcal{W}}^{\prime}$ an MBM class on $M^{[n]}$ and $a,b$ coprime. We write
$\rho^{-1}({\mathcal{W}})={\mathcal{W}}_{\mathrm{K3}}+\lambda\delta,\quad{\mathcal{W}}_{\mathrm{K3}}\in H^{2}(S,{\mathbb{Z}}),\quad\lambda\in{\mathbb{Z}}.$
The equation (13) implies that $\phi_{{\mathcal{U}}}^{-1}(a{\mathcal{W}}_{\mathrm{K3}})$ is an integral class. By [9, Section 3.3], the integrality implies that the pairing
$(H,a{\mathcal{W}}_{\mathrm{K3}})\in{\mathbb{Z}}$
is divisible by
$\frac{16gt^{2}d^{4}}{\mathrm{gcd}\left(16gt^{2}d^{4},4td^{2}k\right)}=g$
where we have used Proposition 2.3(b) in the last equation.
On the other hand, we have
$(H,a{\mathcal{W}}_{\mathrm{K3}})=(H,\rho^{-1}(a{\mathcal{W}}))=(\rho(H),a{\mathcal{W}})=\epsilon({\mathcal{D}},a{\mathcal{W}})+\textup{[integer divisible by $g$]}$
where the last equality uses (11). In particular, we find
$g\,|\,({\mathcal{D}},a{\mathcal{W}})=aC_{1}.$
Combined with Corollary 1.6, this implies
$g\leq a^{2}C_{1}<C_{0}C_{1}$
which contradicts our choice of ${\mathcal{D}}$ in (9). This shows that (13) cannot hold, which proves the proposition. ∎
In conclusion, both (6) and (7) are settled by our choice of the $K3$ surface $S$ and the Mukai vector $v_{0}$; the proof of Theorem 0.3 is now complete. ∎