### v5

# The $D$-equivalence conjecture for hyper-Kähler varieties via hyperholomorphic bundles

Abstract.
We show that birational hyper-Kähler varieties of $K3^{[n]}$-type are derived equivalent, establishing the $D$-equivalence conjecture in these cases. The Fourier–Mukai kernels of our derived equivalences are constructed from projectively hyperholomorphic bundles, following ideas of Markman. Our method also proves a stronger version of the $D$-equivalence conjecture for hyper-Kähler varieties of $K3^{[n]}$-type with Brauer classes.
0. Introduction
Throughout, we work over the complex numbers ${\mathbb{C}}$. We recall that the $D$-equivalence conjecture [5, 20] predicts that birational Calabi–Yau varieties have equivalent bounded derived categories of coherent sheaves.
Conjecture 0.1 ($D$-equivalence conjecture).
If $X,X^{\prime}$ are nonsingular projective birational Calabi–Yau varieties, then there is an equivalence of bounded derived categories
$D^{b}(X)\simeq D^{b}(X^{\prime}).$
The purpose of this paper is to prove Conjecture 0.1 for hyper-Kähler varieties of $K3^{[n]}$-type; these are nonsingular projective varieties deformation equivalent to the Hilbert scheme of $n$ points on a $K3$ surface. More generally, our method reduces the $D$-equivalence conjecture for hyper-Kähler varieties to the construction of certain projectively hyperholomorphic bundles.
Theorem 0.2.
Conjecture 0.1 holds for any hyper-Kähler varieties of $K3^{[n]}$-type.
The $D$-equivalence conjecture has been proven by Bridgeland [6] for Calabi–Yau threefolds. For projective hyper-Kähler fourfolds, the $D$-equivalence conjecture holds by combining the classification results [8, 29] and the case of Mukai flops by Kawamata [20] and Namikawa [25]. However, very few cases of this conjecture are known in dimension $>4$; see [26, 1] for some partial results. Using equivalences obtained from window conditions, Halpern-Leistner [13] proved the $D$-equivalence conjecture for any hyper-Kähler variety which can be realized as a Bridgeland moduli space of stable objects on a (possibly twisted) $K3$ surface. Theorem 0.2 generalizes Halpern-Leistner’s result, but our construction of the derived equivalences is very different. We obtain explicit Fourier–Mukai kernels which rely on the theory of moduli spaces of hyper-Kähler manifolds and hyperholomorphic bundles [28, 23]; this is closer in spirit to the proposal of Huybrechts [17, Section 5.1]. It would be interesting to find connections between the two approaches.
Our method in fact proves the following stronger, twisted version of the $D$-equivalence conjecture involving arbitrary Brauer classes. Let $X\dashrightarrow X^{\prime}$ be a birational transform between hyper-Kähler varieties of ${K3}^{[n]}$-type. It naturally identifies the Brauer groups of $X,X^{\prime}$: any Brauer class $\alpha\in\mathrm{Br}(X)$ induces a Brauer class $\alpha^{\prime}\in\mathrm{Br}(X^{\prime})$.
Theorem 0.3.
Let $X\dashrightarrow X^{\prime}$ be as above, and let $\alpha$ be any Brauer class on $X$. Then there is an equivalence of bounded derived categories of twisted sheaves
$D^{b}(X,\alpha)\simeq D^{b}(X^{\prime},\alpha^{\prime}).$
Theorem 0.3 specializes to Theorem 0.2 by taking $\alpha=0$.
Acknowledgements
We are grateful to Daniel Huybrechts, Zhiyuan Li, Eyal Markman, Alex Perry, and Ziyu Zhang for helpful discussions, and to the anonymous referee for comments and suggestions. D.M. was supported by a Simons Investigator Grant. J.S. was supported by the NSF Grant DMS-2301474 and a Sloan Research Fellowship. R.Z. was supported by the NKRD Program of China No. 2020YFA0713200 and LNMS.
1. Moduli of Hodge isometries
Assume $n\geq 2$. We denote by $\Lambda$ the $K3^{[n]}$-lattice, which is isometric to $H^{2}(X,{\mathbb{Z}})$, equipped with the Beauville–Bogomolov–Fujiki (BBF) form, for any hyper-Kähler manifold $X$ of $K3^{[n]}$-type. [^1] When we say that $X$ is a hyper-Kähler manifold or a manifold of $K3^{[n]}$-type, it means that $X$ is not necessarily projective. In particular, we have a decomposition
$\Lambda=\Lambda_{K3}\oplus{\mathbb{Z}}\delta,\quad\delta^{2}=2-2n$
with $\Lambda_{K3}$ the unimodular $K3$ lattice, so that any vector $\omega\in\Lambda$ can be expressed uniquely as
$\omega=\omega_{K3}+\lambda\delta,\quad\omega_{K3}\in\Lambda_{K3},\quad\lambda\in{\mathbb{Z}}.$
A marking $(X,\eta_{X})$ for a manifold $X$ of $K3^{[n]}$-type is an isometry $\eta_{X}:H^{2}(X,{\mathbb{Z}})\xrightarrow{\simeq}\Lambda$.
1.1. Inseparable pairs
We denote by ${\mathfrak{M}}_{\Lambda}$ the moduli space of marked manifolds $(X,\eta_{X})$ of $K3^{[n]}$-type; it is naturally a non-Hausdorff complex manifold whose non-separation illustrates the complexity of the birational/bimeromorphic geometry of hyper-Kähler varieties/manifolds [14].
We say that a pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is inseparable if they represent inseparable points on the moduli space ${\mathfrak{M}}_{\Lambda}$; as a consequence of the global Torelli theorem, this is equivalent to the condition that $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ share the same period and lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$.
Typical examples of inseparable pairs are given by bimeromorphic transforms. More precisely, a bimeromorphic map $X\dashrightarrow X^{\prime}$ induces a natural identification $H^{2}(X,{\mathbb{Z}})=H^{2}(X^{\prime},{\mathbb{Z}})$ respecting the Hodge structures. A marking $\eta_{X}$ for $X$ then induces a marking $\eta_{X^{\prime}}$ for $X^{\prime}$, and the pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is therefore inseparable. Note that inseparable points are not necessarily induced by bimeromorphic transforms directly. As an example, we consider bimeromorphic $X,X^{\prime}$ as above and assume that
$\rho:H^{2}(X^{\prime},{\mathbb{Z}})\to H^{2}(X,{\mathbb{Z}})$
is a parallel transport respecting the Hodge structures. Then the pair
$(X,\eta_{X}),\quad(X^{\prime},\eta_{X^{\prime}}),\quad\eta_{X^{\prime}}:=\eta_{X}\circ\rho$
is inseparable. By [14] (see also [21, Section 3.1]), every inseparable pair arises this way.
1.2. Hodge isometries
We recall the moduli space of Hodge isometries; this was used by Buskin [9] and Markman [23] to construct algebraic cycles realizing rational Hodge isometries.
For $\phi\in O(\Lambda_{\mathbb{Q}})$, we define ${\mathfrak{M}}_{\phi}$ to be the moduli space of isomorphism classes of quadruples $(X,\eta_{X},Y,\eta_{Y})$ where $(X,\eta_{X}),(Y,\eta_{Y})\in\mathfrak{M}_{\Lambda}$ are the corresponding markings and
$\eta_{Y}^{-1}\circ\phi\circ\eta_{X}:H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$
is a Hodge isometry sending some Kähler class of $X$ to a Kähler class of $Y$. We have the natural forgetful maps
$\Pi_{1}:{\mathfrak{M}}_{\phi}\to{\mathfrak{M}}_{\Lambda},\quad(X,\eta_{X},Y,\eta_{Y})\mapsto(X,\eta_{X}),$
$\Pi_{2}:{\mathfrak{M}}_{\phi}\to{\mathfrak{M}}_{\Lambda},\quad(X,\eta_{X},Y,\eta_{Y})\mapsto(Y,\eta_{Y}).$
Any connected component ${\mathfrak{M}}^{0}_{\phi}$ of ${\mathfrak{M}}_{\phi}$ maps to a connected component of ${\mathfrak{M}}_{\Lambda}$ via $\Pi_{i}$ which we denote by ${\mathfrak{M}}^{0}_{\Lambda}$.
Lemma 1.1 ([23, Lemma 5.7]).
The maps $\Pi_{i}:{\mathfrak{M}}^{0}_{\phi}\to{\mathfrak{M}}^{0}_{\Lambda}$ ($i=1,2$) between connected components are surjective.
Lemma 1.2.
Assume that the point $(X,\eta_{X},Y,\eta_{Y})$ lies in a connected component ${\mathfrak{M}}^{0}_{\phi}$. Assume further that $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ form an inseparable pair such that
(1) $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})\in{\mathfrak{M}}_{\phi}.$
Then $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})$ lies in the same component ${\mathfrak{M}}^{0}_{\phi}$.
Note that (1) is equivalent to the condition that $\eta_{Y}^{-1}\circ\phi\circ\eta_{X^{\prime}}$ sends some Kähler class of $X^{\prime}$ to a Kähler class of $Y$.
Proof.
Both $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$ which we call $\mathfrak{M}^{0}_{\Lambda}$. We first find paths in ${\mathfrak{M}}^{0}_{\Lambda}$ connecting both points to $(X_{0},\eta_{X_{0}})\in{\mathfrak{M}}^{0}_{\Lambda}$ with $\mathrm{Pic}(X_{0})=0$. By Lemma 1.1, we can lift these paths to ${\mathfrak{M}}_{\phi}$, which connect $(X,\eta_{X},Y,\eta_{Y})$ to $(X_{0},\eta_{X_{0}},Y_{0},\eta_{Y_{0}})$, and $(X^{\prime},\eta_{X^{\prime}},Y,\eta_{Y})$ to $(X_{0},\eta_{X_{0}},Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$ respectively. On one hand, by considering the projection $\Pi_{2}$, we know that the two points $(Y_{0},\eta_{Y_{0}}),(Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$ lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$; on the other hand, the Hodge isometry condition ensures that both of them share the same period [23, Lemma 5.4] and they have trivial Picard group. By the global Torelli theorem, we must have $(Y_{0},\eta_{Y_{0}})=(Y^{\prime}_{0},\eta_{Y^{\prime}_{0}})$. This completes the proof. ∎
Suppose we are given a point $(X,\eta_{X},Y,\eta_{Y})$ in ${\mathfrak{M}}_{\phi}$, and Kähler classes $\omega_{X},\omega_{Y}$ on $X,Y$ which are identified via $\eta_{Y}^{-1}\circ\phi\circ\eta_{X}$. Using this data, one can define a diagonal twistor line $\ell\subset{{\mathfrak{M}}}_{\phi}$ which lifts the twistor lines associated to $(X,\omega_{X})$ and $(Y,\omega_{Y})$ on $\mathfrak{M}_{\Lambda}$. A generic diagonal twistor path on ${\mathfrak{M}}_{\phi}$ is given by a chain of diagonal twistor lines such that, at each node of the chain, the associated hyper-Kähler manifolds have trivial Picard group. Generic diagonal twistor paths are used in Theorem 1.3 below to deform certain Fourier–Mukai kernels.
1.3. Brauer groups
Assume that $X$ is a manifold of $K3^{[n]}$-type. Since $X$ has no odd cohomology, the discussion in [11, Section 4.1] yields the following explicit description of the (cohomological) Brauer group:
(2) $\mathrm{Br}(X)=\left(H^{2}(X,{\mathbb{Z}})/\mathrm{Pic}(X)\right)\otimes{\mathbb{Q}}/{\mathbb{Z}}.$
In particular, given a bimeromorphic map $X\dashrightarrow X^{\prime}$ between manifolds of $K3^{[n]}$-type, there is a natural identification
$\mathrm{Br}(X)=\mathrm{Br}(X^{\prime})$
since both $H^{2}(-,{\mathbb{Z}})$ and $\mathrm{Pic}(-)$ are identified for $X$ and $X^{\prime}$. The description (2) also allows us to present a Brauer class in the form
(3) $\left[\frac{\beta}{d}\right]\in\mathrm{Br}(X),\quad\beta\in H^{2}(X,{\mathbb{Z}}),\quad d\in{\mathbb{Z}}_{>0};$
this is referred to as the “$B$-field”.
We note that the cohomology $H^{2}(X,{\mathbb{Z}})$ forms a trivial local system over any connected component of the moduli space ${\mathfrak{M}}^{0}_{\Lambda}$; therefore (3) for a single $X$ presents a Brauer class for any point in the component ${\mathfrak{M}}^{0}_{\Lambda}$ containing $(X,\eta_{X})$.
1.4. Projectively hyperholomorphic bundles
Using the Bridgeland–King–Reid (BKR) correspondence [7], Markman constructed in [23] a class of projectively hyperholomorphic bundles which we recall here. We consider a projective $K3$ surface $S$ with $\mathrm{Pic}(S)={\mathbb{Z}}H$. Assume that $r,s$ are two coprime integers with $r\geq 2$. Assume further that the Mukai vector
$v_{0}:=(r,mH,s)\in H^{*}(S,{\mathbb{Z}})$
is isotropic, i.e. $v_{0}^{2}=0$, and that all stable sheaves on $S$ with Mukai vector $v_{0}$ are stable vector bundles. [^2] In [23], Markman only considered the case $m=1$; here considering large ±mplus-or-$\pm m$ is crucial for our purpose. Using [16, Proposition 2.2] (see also [30, Theorem 2.2]), Markman’s argument works identically in this generality. Let $M$ be the moduli space of such stable vector bundles. Then $M$ is again a $K3$ surface, and the coprime condition of $r,s$ ensures the existence of a universal rank $r$ bundle ${\mathcal{U}}$ on $M\times S$. Conjugating the BKR correspondence, we obtain a vector bundle ${\mathcal{U}}^{[n]}$ on $M^{[n]}\times S^{[n]}$ of rank
$\mathrm{rk}({\mathcal{U}}^{[n]})=n!r^{n};$
see [23, Lemma 7.1]. This vector bundle induces a derived equivalence
(4) $\Phi_{{\mathcal{U}}^{[n]}}:D^{b}(M^{[n]})\xrightarrow{\simeq}D^{b}(S^{[n]}).$
Markman further showed in [23, Section 5.6] that the characteristic class of ${\mathcal{U}}^{[n]}$ induces a Hodge isometry
$\phi_{{\mathcal{U}}^{[n]}}:H^{2}(M^{[n]},{\mathbb{Q}})\to H^{2}(S^{[n]},{\mathbb{Q}}).$
Under the natural identification
(5) $H^{2}(M^{[n]},{\mathbb{Q}})=H^{2}(M,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta,\quad H^{2}(S^{[n]},{\mathbb{Q}})=H^{2}(S,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta,$
this Hodge isometry is of the form
$\phi_{{\mathcal{U}}^{[n]}}=(\phi_{\mathcal{U}},\mathrm{id}),\quad\phi_{\mathcal{U}}:H^{2}(M,{\mathbb{Q}})\to H^{2}(S,{\mathbb{Q}}),$
where $\phi_{\mathcal{U}}$ is the Hodge isometry of $K3$ surfaces induced by ${\mathcal{U}}$; see [23, Corollary 7.3].
The key results, which are summarized in the following theorem, show that the Fourier–Mukai kernel ${\mathcal{U}}^{[n]}$, as a projectively hyperholomorphic bundle, deforms along generic diagonal twistor paths. Moreover, at each point of the path, it induces a (twisted) derived equivalence:
Theorem 1.3 ([23, 19]).
There exist markings $\eta_{M^{[n]}},\eta_{S^{[n]}}$ for the Hilbert schemes $M^{[n]},S^{[n]}$ respectively, which induce $\phi\in O(\Lambda_{\mathbb{Q}})$ via $\phi_{{\mathcal{U}}^{[n]}}$, such that the connected component containing the quadruple
$(M^{[n]},\eta_{M^{[n]}},S^{[n]},\eta_{S^{[n]}})\in{\mathfrak{M}}^{0}_{\phi}$
satisfies the following:
  1. (a)
    For every point $(X,\eta_{X},Y,\eta_{Y})$ lying in the component ${\mathfrak{M}}^{0}_{\phi}$, there exists a twisted vector bundle $({\mathcal{E}},\alpha_{\mathcal{E}})$ on $X\times Y$, which is deformed from ${\mathcal{U}}^{[n]}$ along a generic diagonal twistor path.
  2. (b)
    Using the form (3), the Brauer class in (a) is presented by
    $\alpha_{\mathcal{E}}=\left[-\frac{c_{1}({\mathcal{U}}^{[n]})}{\mathrm{rk}({\mathcal{U}}^{[n]})}\right].$
    Here we view $H^{2}(X\times Y,{\mathbb{Z}})=H^{2}(X,{\mathbb{Z}})\oplus H^{2}(Y,{\mathbb{Z}})$ as a trivial local system over the moduli space ${\mathfrak{M}}^{0}_{\phi}$ via the markings.
  3. (c)
    Further assume that $X,Y$ are varieties. Then the twisted bundle $({\mathcal{E}},\alpha_{\mathcal{E}})$ induces an equivalence of twisted derived categories
    $\Phi_{({\mathcal{E}},\alpha_{\mathcal{E}})}:D^{b}(X,\alpha_{X})\xrightarrow{\simeq}D^{b}(Y,\alpha_{Y}),\quad\alpha_{X}=\left[\frac{a_{X}}{\mathrm{rk}({\mathcal{E}})}\right],\quad\alpha_{Y}=\left[-\frac{a_{Y}}{\mathrm{rk}({\mathcal{E}})}\right],$
    where $a_{X}\in H^{2}(X,{\mathbb{Z}}),a_{Y}\in H^{2}(Y,{\mathbb{Z}})$ are given by
    $c_{1}({\mathcal{U}}^{[n]})=a_{X}+a_{Y}\in H^{2}(X,{\mathbb{Z}})\oplus H^{2}(Y,{\mathbb{Z}}).$
Proof.
(a) was proven in [23, Theorem 8.4]; Markman showed that ${\mathcal{U}}^{[n]}$ on $M^{[n]}\times S^{[n]}$ is projectively slope-stable hyperholomorphic in the sense of [28, 22] which allows him to deform it along diagonal twistor paths to all points in the component ${\mathfrak{M}}^{0}_{\phi}$.
(b) can be obtained by applying Căldăraru’s result [11, Theorem 4.1] along the diagonal twistor paths; see the discussion in [19, Section 2.3].
(c) was proven in [19, Theorem 2.3]. More precisely, the condition that a twisted bundle induces a twisted derived equivalence can be characterized by cohomological properties [10, Theorem 3.2.1]. These properties are preserved along a twistor path due to the fact that the cohomology of slope-polystable hyperholomorphic bundles is invariant under hyper-Kähler rotations [27, Corollary 8.1]. Therefore we ultimately reduce the desired cohomological properties to those for $M^{[n]}\times S^{[n]}$ which are given by the original equivalence (4). ∎
1.5. Birational geometry and MBM classes
The birational geometry of hyper-Kähler varieties is governed by certain integral primitive cohomology classes, called the monodromy birationally minimal (MBM) classes. We refer to [3] for an introduction to these classes. In the following, we summarize some results which are needed in our proof.
Let $X$ be a variety of $K3^{[n]}$-type. We consider its birational Kähler cone $\mathcal{BK}_{X}$ and the positive cone $\mathcal{C}_{X}$:
$\mathcal{BK}_{X}\subset\mathcal{C}_{X}\subset H^{1,1}(X,{\mathbb{R}}).$
The positive cone is convex and admits a wall-and-chamber structure. The closure of the birational Kähler cone within the positive cone is a convex sub-cone [15], which inherits a wall-and-chamber structure. Furthermore, by a result of Amerik–Verbitsky [2], and independently Mongardi [24], all the walls are governed by the MBM classes.
Theorem 1.4 ([24, 2]).
Any wall of $\mathcal{C}_{X}$ is described by a hyperplane of the form
${\mathcal{W}}^{\perp}:=\{\omega\in H^{1,1}(X,{\mathbb{R}}),\,(\omega,{\mathcal{W}})=0\}\subset H^{1,1}(X,{\mathbb{R}})$
with ${\mathcal{W}}$ an algebraic MBM class in $\mathrm{Pic}(X)$. Here the pairing is with respect to the BBF form. Moreover, any chamber in $\mathcal{C}_{X}$ can be realized as the Kähler cone of a birational hyper-Kähler $X^{\prime}$ through a parallel transport $\rho:H^{2}(X^{\prime},{\mathbb{Z}})\to H^{2}(X,{\mathbb{Z}})$ respecting the Hodge structures.
Note that any chamber in $\mathcal{BK}_{X}\subset\mathcal{C}_{X}$ is given by the pullback of the Kähler cone via a birational transform $X\dashrightarrow X^{\prime}$ of hyper-Kähler varieties. By the discussions of Section 1.1, any chamber of $\mathcal{C}_{X}$ corresponds to a marked variety $(X^{\prime},\eta_{X^{\prime}})$ of $K3^{[n]}$-type such that the pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is inseparable.
We also need the following boundedness result, which notably implies that wall-and-chamber structure of $\mathcal{C}_{X}$ is locally polyhedral; see [18, Remark 8.2.3] for a proof of the implication. The boundedness result was essentially obtained by [4], as explained in [2, Section 6.2].
Theorem 1.5 ([4, 2]).
There is a constant $C_{0}>0$, such that for any variety $X$ of $K3^{[n]}$-type and any MBM class ${\mathcal{W}}\in H^{2}(X,{\mathbb{Z}})$ we have
$0<-{\mathcal{W}}^{2}<C_{0}.$
Here the norm is with respect to the BBF form.
For any rational Hodge isometry $\phi:H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$ between varieties of $K3^{[n]}$-type, which sends an MBM class ${\mathcal{W}}_{X}$ on $X$ to a class proportional to an MBM class ${\mathcal{W}}_{Y}$ on $Y$, there exist coprime integers $a,b$ such that
$\phi({\mathcal{W}}_{X})=\frac{a}{b}{\mathcal{W}}_{Y}.$
The following is an immediate consequence of Theorem 1.5.
Corollary 1.6.
For any $X,Y,\phi,{\mathcal{W}}_{X},{\mathcal{W}}_{Y}$ as above, we have
$a^{2}<C_{0},\quad b^{2}<C_{0}.$
Proof.
Since $\phi$ is an isometry, we have
$\frac{a^{2}}{b^{2}}=\frac{{\mathcal{W}}_{X}^{2}}{{\mathcal{W}}_{Y}^{2}}$
By Theorem 1.5, both $-{\mathcal{W}}_{X}^{2}$ and $-{\mathcal{W}}_{Y}^{2}$ are positive integers $<C_{0}$. The corollary follows from the assumption that $a,b$ are coprime. ∎
1.6. Proof strategy
We discuss the strategy of the proof of Theorem 0.3; Theorem 0.2 is then deduced as a special case.
Let $X$ be a variety of $K3^{[n]}$-type. It suffices to prove Theorem 0.3 for a hyper-Kähler birational model $X^{\prime}$ with a birational map $X\dashrightarrow X^{\prime}$ which corresponds to a chamber in $\mathcal{BK}_{X}$ adjacent to the Kähler cone of $X$. By Theorem 1.5, the wall between these two chambers is given by an algebraic MBM class ${\mathcal{W}}\in\mathrm{Pic}(X)$.
Now we choose a $K3$ surface $S$ and a Mukai vector $v_{0}=(r,mH,s)$ as in the beginning of Section 1.4, which yields the Hodge isometry $\phi_{{\mathcal{U}}^{[n]}}$. Associated to these, we have the moduli space of Hodge isometries ${\mathfrak{M}}_{\phi}$, and the component ${\mathfrak{M}}^{0}_{\phi}$ that contains the quadruple $(M^{[n]},\eta_{M^{[n]}},S^{[n]},\eta_{S^{[n]}})$.
For the given birational $X,X^{\prime}$, by Lemma 1.1, we can complete them to a pair of quadruples
$(Y,\eta_{Y},X,\eta_{X}),\quad(Y^{\prime},\eta_{Y^{\prime}},X^{\prime},\eta_{X^{\prime}})\in{\mathfrak{M}}^{0}_{\phi}$
such that the marking $\eta_{X^{\prime}}$ is induced by $\eta_{X}$ via the birational map $X\dashrightarrow X^{\prime}$. [^3] Here we would like $X,X^{\prime}$ to be deformed from $S^{[n]}$ later in Section 2. In particular, the pair $(X,\eta_{X}),(X^{\prime},\eta_{X^{\prime}})$ is inseparable. We note that the pair $(Y,\eta_{Y}),(Y^{\prime},\eta_{Y^{\prime}})$ is also inseparable. This is because they share the same period and lie in the same connected component of ${\mathfrak{M}}_{\Lambda}$. Moreover, by definition, $\phi^{-1}$ sends a Kähler class of $X$ (resp. $X^{\prime}$) to a Kähler class of $Y$ (resp. $Y^{\prime}$). [^4] Here we suppress the markings and still use $\phi$ to denote the Hodge isometry $H^{2}(Y,{\mathbb{Q}})\to H^{2}(X,{\mathbb{Q}})$ for notational convenience. Therefore, if
[…]
there must be a point on the wall separating the Kähler cones of $X,X^{\prime}$ which is sent to the interior of a chamber of the positive cone $\mathcal{C}_{Y}$. In particular, there exists a hyper-Kähler birational model ${Y^{\prime\prime}}$ of $Y$ with a marking $({Y^{\prime\prime}},\eta_{Y^{\prime\prime}})$ such that the pair $(Y,\eta_{Y}),(Y^{\prime\prime},\eta_{Y^{\prime\prime}})$ is inseparable and
$(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X,\eta_{X}),\quad(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X^{\prime},\eta_{X^{\prime}})\in{\mathfrak{M}}_{\phi}.$
Furthermore, by Lemma 1.2, both points lie in the connected component we started with:
$(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X,\eta_{X}),\quad(Y^{\prime\prime},\eta_{Y^{\prime\prime}},X^{\prime},\eta_{X^{\prime}})\in{\mathfrak{M}}^{0}_{\phi}.$
By Theorem 1.3(b, c), we obtain Brauer classes $\alpha_{X},\alpha_{Y^{\prime\prime}}$ on $X,Y^{\prime\prime}$ respectively, such that
$D^{b}(Y^{\prime\prime},\alpha_{Y^{\prime\prime}})\simeq D^{b}(X,\alpha_{X}),\quad D^{b}(Y^{\prime\prime},\alpha_{Y^{\prime\prime}})\simeq D^{b}(X^{\prime},\alpha_{X^{\prime}}).$
Here the Brauer classes $\alpha_{X},\alpha_{Y^{\prime\prime}}$ only depend on the markings $(X,\eta_{X}),(Y^{\prime\prime},\eta_{Y^{\prime\prime}})$ respectively, and the Brauer class $\alpha_{X^{\prime}}$ is induced by $\alpha_{X}$. Combining both equivalences yields
$D^{b}(X,\alpha_{X})\simeq D^{b}(X^{\prime},\alpha_{X^{\prime}})$
whose Fourier–Mukai kernel is the composition of two (twisted) hyperholomorphic bundles.
In the next section, we show that for any pair $X,X^{\prime}$ as above with a Brauer class $\alpha\in\mathrm{Br}(X)$ and an algebraic MBM class ${\mathcal{W}}\in\mathrm{Pic}(X)$, a careful choice of the $K3$ surface $S$ and the Mukai vector $v_{0}=(r,mH,s)$ as in Section 1.4 can simultaneously ensure that the condition (6) holds and the induced Brauer class is as desired:
(7) $\alpha_{X}=\alpha.$
This completes the proof of Theorem 0.3.
Remark 1.7.
For a general birational transform $X\dashrightarrow X^{\prime}$ of varieties of $K3^{[n]}$-type, which does not correspond to adjacent chambers in the birational Kähler cone $\mathcal{BK}_{X}$, our proof realizes the derived equivalence
$D^{b}(X,\alpha)\simeq D^{b}(X^{\prime},\alpha^{\prime})$
via two sequences of varieties $X_{1},\ldots,X_{t-1}$ and $Y_{1},\ldots,Y_{t}$, with each $X_{i}$ birational to $X,X^{\prime}$, such that
(8) $D^{b}(X,\alpha)\simeq D^{b}(Y_{1},\alpha_{Y_{1}})\simeq D^{b}(X_{1},\alpha_{X_{1}})\simeq D^{b}(Y_{2},\alpha_{Y_{2}})\simeq\cdots\\ \simeq D^{b}(Y_{t-1},\alpha_{Y_{t-1}})\simeq D^{b}(X_{t-1},\alpha_{X_{t-1}})\simeq D^{b}(Y_{t},\alpha_{Y_{t}})\simeq D^{b}(X^{\prime},\alpha^{\prime}).$
Each of the derived equivalences in (8) is induced by a (twisted) hyperholomorphic bundle.
2. Proof of Theorem 0.3
From now on, we fix a variety $X$ of $K3^{[n]}$-type, a Brauer class $\alpha\in\mathrm{Br}(X)$, and an algebraic MBM class ${\mathcal{W}}\in\mathrm{Pic}(X)$ as in Section 1.6. In particular, the variety $X$ has Picard rank 2absent2$\geq$ 2≥ 2.555Theorem 0.3 is automatically true if $X$ has Picard rank $1$, since any birational transform $X\dashrightarrow X^{\prime}$ is necessarily an isomorphism. Using (2) and (3), we present the Brauer class $\alpha$ by a class in the rational transcendental lattice $T(X)_{\mathbb{Q}}\subset H^{2}(X,{\mathbb{Q}})$:
$\alpha=\left[-\frac{{\mathcal{B}}}{d}\right],\quad{\mathcal{B}}\in T(X),\quad d\in{\mathbb{Z}}_{>0}.$
Up to adjusting $-\frac{{\mathcal{B}}}{d}$ by an integral class in $T(X)\subset H^{2}(X,{\mathbb{Z}})$, we may further assume that the class ${\mathcal{B}}$ satisfies
${\mathcal{B}}^{2}=2e>0.$
2.1. Divisor classes
Recall that the divisibility $\mathrm{div}(\omega)$ of a class $\omega\in H^{2}(X,{\mathbb{Z}})$ is the positive generator of the subgroup
$\{(\omega,\mu),\,\mu\in H^{2}(X,{\mathbb{Z}})\}\subset{\mathbb{Z}}$
Lemma 2.1.
There exists a class ${\mathcal{A}}\in\mathrm{Pic}(X)$ such that
$({\mathcal{A}},{\mathcal{W}})\neq 0,\quad\mathrm{div}({\mathcal{A}})=1.$
Proof.
We pick a marking identifying $H^{2}(X,{\mathbb{Z}})$ with a $K3^{[n]}$-lattice $\Lambda=\Lambda_{K3}\oplus{\mathbb{Z}}\delta$. For any $g\in O(\Lambda)$, since $g(\delta)^{\perp}$ is a unimodular $K3$-lattice, any primitive vector $\omega\in g(\delta)^{\perp}\subset\Lambda$ satisfies $\mathrm{div}(\omega)=1.$ We would like to choose $g$ so that there exists ${\mathcal{A}}\in g(\delta)^{\perp}\cap\mathrm{Pic}(X)$ satisfying $({\mathcal{A}},{\mathcal{W}})\neq 0$. In other words, we want
$g(\delta)^{\perp}\cap\mathrm{Pic}(X)\neq{\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X).$
If we base change to $\mathbb{C}$, the set of $g\in O(\Lambda)_{\mathbb{C}}$ such that
$g(\delta)^{\perp}\cap\mathrm{Pic}(X)_{\mathbb{C}}\neq{\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X)_{\mathbb{C}}$
is open in the Zariski topology. Furthermore, it is nonempty since $X$ has Picard rank $\geq 2$. Since $O(\Lambda)$ is Zariski-dense in $O(\Lambda)_{\mathbb{C}}$, we can find $g\in O(\Lambda)$ satisfying this condition as well. ∎
Up to replacing ${\mathcal{A}}$ by $-{\mathcal{A}}$, we may assume
$C_{1}:=({\mathcal{A}},{\mathcal{W}})>0$
which we fix from now on.
Proposition 2.2.
For any $N>0$, there exists a class ${\mathcal{D}}\in\mathrm{Pic}(X)$ of divisibility $1$, satisfying
${\mathcal{D}}^{2}>N,\quad({\mathcal{D}},{\mathcal{W}})=C_{1}.$
Proof.
Since $X$ has Picard rank $\geq 2$, we have ${\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X)\neq 0$. Pick an element
$\omega\in{\mathcal{W}}^{\perp}\cap\mathrm{Pic}(X),\quad\omega^{2}>0.$
Then for large enough $t\in{\mathbb{Z}}_{>0}$, we have
$({\mathcal{A}}+t\omega,{\mathcal{W}})=C_{1},\quad({\mathcal{A}}+t\omega)^{2}>N.$
It suffices to show that there exist infinitely many choices of $t\in{\mathbb{Z}}_{>0}$ satisfying
$\mathrm{div}({\mathcal{A}}+t\omega)=1.$
We pick an integral class $\mu\in H^{2}(X,{\mathbb{Z}})$ such that
$({\mathcal{A}},\mu)=1,\quad(\omega,\mu)\neq 0;$
then we pick another integral class $\nu\in H^{2}(X,{\mathbb{Z}})$ such that
$({\mathcal{A}},\nu)=0,\quad(\omega,\nu)\neq 0.$
By Dirichlet’s theorem on primes in arithmetic progressions, there exists a sufficiently large $t\in{\mathbb{Z}}_{>0}$ such that the absolute value of $1+t(\omega,\mu)$ is a prime number. We claim that for such a choice of $t$, the class ${\mathcal{A}}+t\omega$ must have divisibility $1$. This follows immediately from the observation that
$\mathrm{div}({\mathcal{A}}+t\omega)\mid 1+t(\omega,\mu),\quad\mathrm{div}({\mathcal{A}}+t\omega)\mid(\omega,\nu).\qed$
2.2. Mukai vectors
We construct the $K3$ surface $S$ and the Mukai vector $v_{0}$ of Section 1.6.
By Proposition 2.2, we can find ${\mathcal{D}}\in\mathrm{Pic}(X)$ with
(9) $\mathrm{div}({\mathcal{D}})=1,\quad({\mathcal{D}},{\mathcal{W}})=C_{1}>0,\quad{\mathcal{D}}^{2}=2g>2C_{0}C_{1},$
where $C_{0}$ is the constant in Theorem 1.5. Repeating the same argument as in Proposition 2.2, we also find $t\in{\mathbb{Z}}_{>0}$ such that
$\mathrm{div}({\mathcal{D}}+4gtd{\mathcal{B}})=1.$
Let $(S,H)$ be a primitively polarized $K3$ surface of Picard rank $1$ of degree
$H^{2}=2g\left(1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)>0.$
We observe that both classes
$H-2gtd^{2}\delta\in\mathrm{H}^{2}(S^{[n]},{\mathbb{Z}}),\quad{\mathcal{D}}+4gtd{\mathcal{B}}\in H^{2}(X,{\mathbb{Z}})$
are of divisibility 1 and have the same norm, where we have used that $({\mathcal{D}},{\mathcal{B}})=0$ since ${\mathcal{B}}$ is transcendental. Therefore, by [12, Example 3.8] and [21, Theorem 9.8], there is a parallel transport
$\rho:H^{2}(S^{[n]},{\mathbb{Z}})\to H^{2}(X,{\mathbb{Z}})$
satisfying
(10) $\rho(H-2gtd^{2}\delta)=\epsilon({\mathcal{D}}+4gtd{\mathcal{B}}),$
where $\epsilon=\pm 1$ is a sign determined by the orientation.
We now consider the Mukai vector
[…]
which clearly satisfies
$\mathrm{gcd}\left(16gt^{2}d^{4},1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)=1,\quad 16gt^{2}d^{4}\geq 2,\quad v_{0}^{2}=0.$
Since $g>1$, we have
$14gtd^{2}\nmid\frac{H^{2}}{2}+1=g\left(1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)+$ 1
which by [30, Lemma 1.2] implies that all stable sheaves on $S$ with Mukai vector $v_{0}$ are stable vector bundles. Therefore, the moduli space $M$ of stable vector bundles on $S$ with Mukai vector $v_{0}$ is a $K3$ surface of Picard rank $1$ with a universal bundle ${\mathcal{U}}$ on $M\times S$ which we fix from now on. Also fixed are the markings $\eta_{M^{[n]}},\eta_{S^{[n]}}$ as in Theorem 1.3, as well as the induced marking
$\eta_{X}:=\eta_{S^{[n]}}\circ\rho^{-1}:H^{2}(X,{\mathbb{Z}})\xrightarrow{\simeq}\Lambda.$
Proposition 2.3.
Let $S,M,{\mathcal{U}}$ be as above.
  1. (a)
    The primitive polarization $\widehat{H}$ of $M$ satisfies $\widehat{H}^{2}=H^{2}$.
  2. (b)
    Let $s\in S$ be a point. Assume that the vector bundle ${\mathcal{U}}|_{s}$ has Mukai vector
    $\widehat{v}_{0}=(16gt^{2}d^{4},k\widehat{H},\widehat{s})\in H^{*}(M,{\mathbb{Z}}).$
    Then we have
    $\mathrm{gcd}(16gt^{2}d^{4},k)=4td^{2}.$
Proof.
(a) follows from [30, Appendix A]. For (b), we note that [30, Theorem 2.2] implies that the Mukai vector $\widehat{v}_{0}$ is primitive with $\widehat{v}_{0}^{2}=0$. Using (a), we deduce that
$\widehat{s}=\left(\frac{k}{4td^{2}}\right)^{2}\left(1+4gt^{2}d^{4}(n-1)+16gt^{2}d^{2}e\right)\in{\mathbb{Z}}.$
Therefore, we have that $k$ is divisible by $4td^{2}$, which shows
$4td^{2}\mid\mathrm{gcd}(16gt^{2}d^{4},k).$
On the other hand, if $\frac{k}{4td^{2}}$ is not coprime to $16gt^{2}d^{4}$, the Mukai vector $\widehat{v}_{0}$ is divisible by their common factor. This contradicts the fact that $\widehat{v}_{0}$ is primitive. ∎
2.3. End of proof
We complete the proof using the $K3$ surface $S$, the Mukai vector $v_{0}$, and the universal bundle ${\mathcal{U}}$ constructed in the last section. This gives the vector bundle ${\mathcal{U}}^{[n]}$ on $M^{[n]}\times S^{[n]}$. We write
$c_{1}({\mathcal{U}}^{[n]})=a_{M^{[n]}}+a_{S^{[n]}}\in H^{2}(M^{[n]},{\mathbb{Z}})\oplus H^{2}(S^{[n]},{\mathbb{Z}})$
with
$a_{M^{[n]}}\in H^{2}(M^{[n]},{\mathbb{Z}}),\quad a_{S^{[n]}}\in H^{2}(S^{[n]},{\mathbb{Z}}).$
Recall the natural identification
(11) $H^{2}(S^{[n]},{\mathbb{Z}})=H^{2}(S,{\mathbb{Z}})\oplus{\mathbb{Z}}\delta.$
By [23, Equation (7.11)], we can present the class $a_{S^{[n]}}$ using (11):
$a_{S^{[n]}}=\mathrm{rk}({\mathcal{U}}^{[n]})\cdot\left(\frac{\epsilon\cdot 4td^{2}H}{16gt^{2}d^{4}}-\frac{\delta}{2}\right)\in H^{2}(S^{[n]},{\mathbb{Z}}).$
Via the parallel transport $\rho$ and (10), we obtain
$\rho\left(\frac{a_{S^{[n]}}}{\mathrm{rk}({\mathcal{U}}^{[n]})}\right)$ $=\rho\left(\frac{\epsilon\cdot H}{4gtd^{2}}-\frac{\delta}{2}\right)$
$=\rho\left(\epsilon\left(\frac{H}{4gtd^{2}}-\frac{\delta}{2}\right)+(\epsilon-1)\frac{\delta}{2}\right)$
$=\frac{\epsilon\cdot\rho(H-2gtd^{2}\delta)}{4gtd^{2}}+\frac{(\epsilon-1)}{2}\rho(\delta)$
$=\frac{{\mathcal{D}}+4gtd{\mathcal{B}}}{4gtd^{2}}+\frac{(\epsilon-1)}{2}\rho(\delta)$
$=\frac{{\mathcal{B}}}{d}+\textup{[class in $\mathrm{Pic}(X)_{\mathbb{Q}}$]}+\textup{[class in $H^{2}(X,{\mathbb{Z}})$]}\in H^{2}(X,{\mathbb{Q}}).$
Hence, by Theorem 1.3(c), we have
$\alpha_{X}=\left[-\rho\left(\frac{a_{S^{[n]}}}{\mathrm{rk}({\mathcal{U}}^{[n]})}\right)\right]=\left[-\frac{{\mathcal{B}}}{d}\right]=\alpha.$
To complete the proof, it remains to address (6). This is given by the following proposition.
Proposition 2.4.
Let ${\mathfrak{M}}_{\phi}^{0}$ be the connected component of the moduli space of Hodge isometries constructed from $S,M,{\mathcal{U}}$ as above. For any quadruple
$(Y,\eta_{Y},X,\eta_{X})\in{\mathfrak{M}}_{\phi}^{0}$
with $X,{\mathcal{W}}$ fixed as above, the class $\phi^{-1}({\mathcal{W}})\in H^{2}(Y,{\mathbb{Q}})$ is not proportional to any MBM class on $Y$. Here we suppress the markings and still use $\phi$ to denote the Hodge isometry $H^{2}(Y,{\mathbb{Q}})\to H^{2}(X,{\mathbb{Q}})$ for notational convenience.
Proof.
The main idea of the argument is that, for our choice of the Mukai vector $v_{0}$, by a calculation of Buskin [9], the rational Hodge isometry $\phi^{-1}$ is conjugate to a reflection by a vector of large norm. By Corollary 1.6, we then show that it cannot send ${\mathcal{W}}$ to a class proportional to an MBM class.
The details are as follows. Since the MBM classes are deformation invariant, we only need to treat the Hodge isometry
$\phi_{{\mathcal{U}}^{[n]}}:H^{2}(M^{[n]},{\mathbb{Q}})\to H^{2}(S^{[n]},{\mathbb{Q}})$
which can be further simplified under the identification (5):
$(\phi_{\mathcal{U}},\mathrm{id}):H^{2}(M,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta\to H^{2}(S,{\mathbb{Q}})\oplus{\mathbb{Q}}\delta.$
Assume that
(12) $\phi_{{\mathcal{U}}^{[n]}}^{-1}(\rho^{-1}({\mathcal{W}}))=\frac{b}{a}{\mathcal{W}}^{\prime}$
with ${\mathcal{W}}^{\prime}$ an MBM class on $M^{[n]}$ and $a,b$ coprime. We write
$\rho^{-1}({\mathcal{W}})={\mathcal{W}}_{K3}+\lambda\delta,\quad{\mathcal{W}}_{K3}\in H^{2}(S,{\mathbb{Z}}),\quad\lambda\in{\mathbb{Z}}.$
The equation (12) implies that $\phi_{{\mathcal{U}}}^{-1}(a{\mathcal{W}}_{K3})$ is an integral class. By the formula right before [9, Conclusion 3.8], the integrality forces the pairing
$(H,a{\mathcal{W}}_{K3})\in{\mathbb{Z}}$
to be divisible by
$\frac{16gt^{2}d^{4}}{\mathrm{gcd}\left(16gt^{2}d^{4},4td^{2}k\right)}=g,$
where we have used Proposition 2.3(b) in the last equation.
On the other hand, we have
$(H,a{\mathcal{W}}_{K3})=(H,\rho^{-1}(a{\mathcal{W}}))=(\rho(H),a{\mathcal{W}})=\epsilon({\mathcal{D}},a{\mathcal{W}})+\textup{[integer divisible by $g$]},$
where the last equality uses (10). In particular, we find
$g\mid({\mathcal{D}},a{\mathcal{W}})=aC_{1}.$
Combined with Corollary 1.6, this implies
$g\leq a^{2}C_{1}<C_{0}C_{1}$
which contradicts our choice of ${\mathcal{D}}$ in (9). This shows that (12) cannot hold, which proves the proposition. ∎
In conclusion, both (6) and (7) are settled by our choice of the $K3$ surface $S$ and the Mukai vector $v_{0}$; the proof of Theorem 0.3 is now complete. ∎